### v3

# Tsirelson bounds for quantum correlations with indefinite causal order

Abstract
Quantum theory is in principle compatible with processes that violate causal inequalities, an analogue of Bell inequalities that constrain the correlations observed by sets of parties operating in a definite causal order. Since the introduction of causal inequalities, determining their maximum quantum violation, analogue to Tsirelson’s bound for Bell inequalities, has remained an open problem. Here we provide a general method for bounding the violation of arbitrary causal inequalities, establishing limits to the correlations achievable by arbitrary local experiments and by arbitrary quantum processes with indefinite causal order. We prove that the maximum violation is generally smaller than the algebraic maximum of the corresponding correlation, and determine Tsirelson-like bounds for a class of causal inequalities including some of the most paradigmatic examples. Our results motivate a search for physical principles characterizing the boundary of the set of quantum correlations with indefinite causal order.
I Introduction
Traditional formulations of physics generally assume that events take place in a well-defined causal order. On the other hand, it has been observed that quantum theory is in principle compatible with scenarios where the order of events is indefinite Chiribella et al. (2009a); Oreshkov et al. (2012); Chiribella et al. (2013); Brukner (2014), a phenomenon that has potential applications to quantum information Chiribella (2012); Araújo et al. (2014); Guérin et al. (2016); Ebler et al. (2018); Zhao et al. (2020); Felce and Vedral (2020); Gao et al. (2023); Zhu et al. (2023) and is expected to play an integral role in a future theory of quantum gravity Hardy (2007). The possibility of indefinite causal order gives rise to an analogue of quantum nonlocality, originally shown by Oreshkov, Costa, and Brukner (OCB) Oreshkov et al. (2012), who developed a framework for studying the most general correlations arising from quantum experiments performed in a set of local laboratories. In this scenario, the assumption that the experiments are performed in a definite causal order implies an analogue of Bell inequalities, called causal inequalities. Strikingly, OCB found that the validity of quantum theory in the local laboratories is compatible with correlations that violate a causal inequality, now called the OCB inequality.
Over the past decade, quantum violations of causal inequalities were found in a variety of scenarios Oreshkov and Giarmatzi (2016); Branciard et al. (2015); Baumeler et al. (2014); Feix et al. (2016); Abbott et al. (2016), also including scenarios achievable with known physics using time-delocalized quantum systems Wechs et al. (2023). These findings raise a fundamental question: what is the maximum violation of causal inequalities that is logically consistent with the validity of quantum theory in a set of local laboratories? Can it reach the maximum algebraic value of the corresponding correlations? Answering these questions is important for understanding the extent to which quantum mechanics is compatible with indefinite causal order. In particular, if the maximum quantum-compatible violation turns out to be smaller than the maximum algebraic value, then the knowledge of the maximum quantum value provides the basis for a search for physical principles explaining why quantum theory is in principle compatible with some amount of causal indefiniteness, but not with the maximal amount. Analogues of these questions have been extensively studied in the context of Bell inequalities Cirel’son (1980); Popescu and Rohrlich (1994); Van Dam (1999); Brassard et al. (2006); Navascués et al. (2007); Brunner and Skrzypczyk (2009); Linden et al. (2007); Pawłowski et al. (2009); Navascués and Wunderlich (2010); Fritz et al. (2013), where they provided valuable insights into the physical and information-theoretic principles underlying quantum theory. In the paradigmatic case of the Clauser-Horne-Shimony-Holt (CHSH) inequality Clauser et al. (1969), the maximum quantum violation was identified by Tsirelson Cirel’son (1980) and is now known as the Tsirelson bound.
In stark contrast with the Bell scenario, little is known about the boundary of the set of quantum correlations with indefinite causal order. For the OCB inequality, an upper bound on the violations achieved by a restricted class of local experiments was derived in Ref. Brukner (2015), but whether larger violations could be achieved by more general experiments remained as an open question. For another causal inequality, associated to the Guess-Your-Neighbor’s Input game Almeida et al. (2010), it was recently proven that the maximum violation has to be strictly smaller than the algebraic maximum Kunjwal and Oreshkov (2023). In general, however, no upper bound other than the algebraic maximum has been found for any causal inequality so far.
Here we develop a general method for bounding the maximum violation of arbitrary causal inequalities by quantum processes with indefinite causal order. As an application of the general method, we establish the analogue of Tsirelson’s bound for the OCB inequality and for a class of causal inequalities, which we name single-trigger inequalities. We then ask whether there exist alternative physical theories that allow for the maximum algebraic violation of causal inequalities. We answer the question in the affirmative, by showing two variants of classical and quantum theory that are in principle compatible with correlations that reach the maximum algebraic value of all causal inequalities with two parties, with up to three settings per party. Overall, our findings open up a search for physical principles determining the boundaries of the set of quantum correlations with indefinite order, potentially leading to a new axiomatization of quantum theory in that does not presuppose a pre-defined causal structure.
II Results
Single-trigger causal inequalities. Here we introduce a special class of causal inequalities that provide the foundation of our method. For the inequalities in this class, the maximum violation can be determined explicitly by a semidefinite program, which in turn can be used to provide upper bounds to the violation of arbitrary causal inequalities.
In the framework of causal inequalities Oreshkov et al. (2012); Oreshkov and Giarmatzi (2016); Abbott et al. (2016), a set of parties operate in different regions of spacetime, performing local operations in their laboratories. The interaction between the parties’ laboratories and the outside world takes place only at specific moments: in the simplest presentation of the framework, the $i$-th laboratory is assumed to be shielded from the outside world at all times, except for two moments $t_{i}$ and $t_{i}^{\prime}\geq t_{i}$ when a shutter is opened, allowing physical systems to enter and exit the laboratory, respectively Oreshkov et al. (2012).
In the time between $t_{i}$ and $t_{i}^{\prime}$ the $i$-th party performs an experiment, obtaining an outcome. We denote by $x_{i}$ ($a_{i}$) the setting (outcome) of the experiment performed by the $i$-th party, and by $\vec{x}=(x_{1},\dots,x_{N})$ ($\vec{a}=(a_{1},\dots,a_{N})$) the vector of all parties’ settings (outcomes). An $N$-partite correlation function is an expression of the form
$\mathcal{I}=\sum_{\vec{a},\vec{x}}\,\alpha_{\vec{a},\vec{x}}~{}p(\vec{a}\,|\,\vec{x})\,,$ (1)
where each $\alpha_{\vec{a},\vec{x}}$ is a real coefficient and $p(\vec{a}\,|\,\vec{x})$ is the conditional probability distribution of the outcomes given the settings.
When the parties operate in a definite causal order, the probability distribution $p(\vec{a}\,|\,\vec{x})$ is subject to a set of linear constraints Oreshkov et al. (2012); Oreshkov and Giarmatzi (2016); Abbott et al. (2016). In the case of $N=2$ parties, Alice and Bob, the constraints have a simple expression: if Alice’s experiment precedes Bob’s experiment, then the probability distribution of Alice’s outcomes must be independent of Bob’s settings, namely $p_{A}(a_{1}\,|\,x_{1},x_{2})=p_{A}(a_{1}\,|\,x_{1},x_{2}^{\prime})$, $\forall x_{2},x_{2}^{\prime}$, with $p_{A}(a_{1}\,|\,x_{1},x_{2}):=\sum_{a_{2}}\,p(a_{1},a_{2}\,|\,x_{1},x_{2})$. Vice-versa, if Bob’s experiment precedes Alice’s experiment, then Bob’s outcomes must be independent of Alice’s settings, namely $p_{B}(a_{2}\,|\,x_{1},x_{2})=p_{B}(a_{2}\,|\,x_{1}^{\prime},x_{2})\,,\forall x_{1},x_{1}^{\prime}$, with $p_{B}(a_{2}\,|\,x_{1},x_{2}):=\sum_{a_{1}}\,p(a_{1},a_{2}\,|\,x_{1},x_{2})$. A probability distribution $p(a_{1},a_{2}\,|\,x_{1},x_{2})$ is called causal if it is a random mixture of probability distributions corresponding to scenarios in which either Alice’s experiment precedes Bob’s or Bob’s experiment precedes Alice’s Oreshkov et al. (2012). In the multipartite case, causal probability distributions can arise in a more general way, by dynamically controlling the order of some of the parties based on outcomes obtained by some of the other parties Oreshkov and Giarmatzi (2016); Abbott et al. (2016).
A causal inequality is an upper bound on the correlations achievable by causal probability distributions; explicitly, it is a bound of the form $\mathcal{I}^{\rm causal}\leq\beta$, where $\mathcal{I}^{\rm causal}$ is the maximum correlation achieved by causal probability distributions and $\beta\in\mathbb{R}$ is some constant. The first example of a causal inequality was introduced by OCB Oreshkov et al. (2012), who showed that quantum theory is in principle compatible with its violation.
In general, the violation of causal inequalities takes place when the experiments performed by the different parties are connected in an indefinite order. The connections are implemented by suitable processes, which are in principle compatible with the validity of quantum theory in the parties’ local laboratories Oreshkov et al. (2012); Chiribella et al. (2013); Araújo et al. (2015); Chiribella and Ebler (2016); Bisio and Perinotti (2019). In the $N$-partite setting, a process of this kind is represented by a linear map $\mathcal{S}$ that transforms the parties’ local operations into the conditional probability distribution
$p(\vec{a}\,|\,\vec{x})=\mathcal{S}\left(\mathcal{M}^{(1)}_{a_{1}\,|\,x_{1}},\cdots,\mathcal{M}^{(N)}_{a_{N}\,|\,x_{N}}\right)\,,$ (2)
where, for every $i\in\{1,\dots,N\}$, $\mathcal{M}^{(i)}_{a_{i}\,|\,x_{i}}$ is the quantum operation occurring in the $i$-th laboratory when an experiment with setting $x_{i}$ produces the outcome $a_{i}$. The positivity and normalization of the probability distributions $p(\vec{a}\,|\,\vec{x})$ for every possible set of experiments performed in the parties’ laboratories place constraints on the admissible maps $\mathcal{S}$ (see Appendix A for details.) The set of maps $\mathcal{S}$ satisfying these constrains contains all possible processes with definite causal order, as well as another type of processes, hereafter referred as quantum processes with indefinite causal order (ICO).
Given a correlation function $\mathcal{I}$, a fundamental problem is to determine the maximum of $\mathcal{I}$ over all possible probability distributions (2) generated by local quantum experiments and by quantum processes with ICO. In the following, we will denote the maximum by $\mathcal{I}^{\rm ICO}$, and call it the ICO bound. The ICO bound is an analogue of the Tsirelson bound Cirel’son (1980) for causal inequalities: the value of $\mathcal{I}^{\rm ICO}$ determines whether quantum mechanics allows for a violation of the causal inequality for the correlation $\mathcal{I}$, and, in the affirmative case, it provides the maximum violation in principle compatible with the validity of quantum theory in the parties’ local laboratories. It is worth noting that, while the physical realization of general ICO processes is currently an open problem, an important subclass of ICO processes can be realized using time-delocalized quantum systems Oreshkov (2019) and some processes in this subclass have been shown to violate causal inequalities Wechs et al. (2023). Since the ICO bound limits the violations achievable by arbitrary ICO processes, in particular it provides an upper bound on the maximum violations arising from time-delocalized quantum systems, thereby providing a non-trivial constraint on the observable consequences of delocalization in time.
Finding the maximum of a correlation over all local experiments and over all ICO processes is a difficult optimization problem, which in principle requires a maximization over quantum systems of arbitrary dimensions, as in the case of Bell correlations. Compared to the maximization of Bell correlations, however, the maximization of ICO correlations appears to be a harder problem, and the exact value of the ICO bound has remained unknown until now for all causal inequalities, except those that have been shown to be violated to their algebraic maximum Baumeler et al. (2014); Baumeler and Wolf (2016); Abbott et al. (2016).
We now introduce a class of $N$-partite correlations for which the maximum quantum violation can be computed explicitly. These correlations, called single-trigger correlations, can be thought as the score achieved by the parties in a game where each party is asked a question and the payoff depends on the party’s answer only if the party receives a specific question, called the “trigger.” The precise definition is as follows: an $N$-partite correlation (1) is single-trigger if for every party $i$ there exists one and only one setting $\xi_{i}$ (the “trigger”) such that $\alpha_{\vec{a},\vec{x}}$ depends on $a_{i}$ only if $x_{i}=\xi_{i}$.
[…]
Crucially, every correlation $\mathcal{I}$ can be decomposed into a sum of single-trigger correlations, and the number of non-zero terms in the sum is at most equal to the total number of settings, namely $n_{1}\cdot n_{2}\cdots n_{N}$, where $n_{i}$ is the number of settings for the $i$-th party. Indeed, the existence of a decomposition with this number of terms is immediate, as one can always write the correlation $\mathcal{I}=\sum_{\vec{a},\vec{x}}\,\alpha_{\vec{a},\vec{x}}\,p(\vec{a}\,|\,\vec{x})$ as $\mathcal{I}=\sum_{\vec{\xi}}\mathcal{I}_{\vec{\xi}}$, where $\mathcal{I}_{\vec{\xi}}$ is the single-trigger correlation with coefficients $\alpha^{\vec{\xi}}_{\vec{a},\vec{x}}:=\delta_{\vec{x},\vec{\xi}}\,\alpha_{\vec{\xi},\vec{a}}$. In general, there can be other decompositions with a smaller number of non-zero terms.
In the following, the causal inequalities associated to single-trigger correlations will be called “single-trigger causal inequalities.” For these inequalities, we will provide an explicit expression of the ICO bound. Since every possible correlation can be decomposed into a sum of single-trigger correlations, this bound will imply general upper bounds on the violation of arbitrary causal inequalities.
Maximum quantum violation of single-trigger causal inequalities. We now provide the central result of the paper: the ICO bound for single-trigger causal inequalities. A crucial feature of our result is that it reduces the problem of calculating the ICO bound to a semidefinite program (SDP) on a space of fixed dimension, depending only on the number of parties, settings, and outcomes.
In addition to the value of the ICO bound, our result provides a canonical choice of local experiments that achieve the bound. The canonical choice for the $i$-th party is to use as input a quantum system of dimension $m_{i}$, equal to the number of possible outcomes, and to append to it an auxiliary system of dimension $n_{i}$, equal to the number of possible settings. In other words, the canonical strategy uses an input system with Hilbert space $\mathcal{H}^{(i)}_{\rm in}=\mathbb{C}^{m_{i}}$ and an output system with Hilbert space $\mathcal{H}^{(i)}_{\rm out}=\mathcal{H}^{(i)}_{\rm in}\otimes\mathcal{H}^{(i)}_{\rm aux}$ with $\mathcal{H}^{(i)}_{\rm aux}=\mathbb{C}^{n_{i}}$. The auxiliary system is used by the $i$-th party to communicate their setting to the outside world: explicitly, the value of the setting will be encoded in the pure state $|x_{i}\rangle\langle x_{i}|$, where $\{|x_{i}\rangle\}_{x_{i}=1}^{n_{i}}$ is a fixed orthonormal basis for $\mathcal{H}^{(i)}_{\rm aux}$. If the setting is equal to the trigger, then the party will measure the input system on the canonical basis for $\mathcal{H}_{\rm in}^{(i)}$. If the setting is not equal to the trigger, the party will not perform any measurement and will just output a random outcome. Mathematically, this strategy is described by quantum operations of the form $\mathcal{M}^{(i)*}_{a_{i}\,|\,x_{i}}=\mathcal{N}^{(i)*}_{a_{i}\,|\,x_{i}}\otimes|x_{i}\rangle\langle x_{i}|$, where $\mathcal{N}^{(i)*}_{a_{i}\,|\,x_{i}}$ is the quantum operation defined by $\mathcal{N}^{(i)*}_{a_{i}\,|\,x_{i}}(\rho):=\langle a_{i}|\rho|a_{i}\rangle\,|a_{i}\rangle\langle a_{i}|$ if $x_{i}=\xi_{i}$ and $\mathcal{N}^{(i)*}_{a_{i}\,|\,x_{i}}(\rho):=\rho/m_{i}$ if $x_{i}\not=\xi_{i}$, for every density matrix $\rho$.
The value of the ICO bound is expressed in terms of the canonical strategy. To this purpose, we define the linear map $\mathcal{M}^{*}_{\mathcal{I}}=\sum_{\vec{a},\vec{x}}\,\alpha_{\vec{a},\vec{x}}\,\mathcal{M}^{*}_{\vec{a}\,|\,\vec{x}}$, where $\alpha_{\vec{a},\vec{x}}$ are the coefficients of the correlation $\mathcal{I}$ [cf. Eq. (1)] and $\mathcal{M}^{*}_{\vec{a}\,|\,\vec{x}}:=\mathcal{M}^{(1)*}_{a_{1}\,|\,x_{1}}\otimes\mathcal{M}^{(2)*}_{a_{2}\,|\,x_{2}}\otimes\cdots\otimes\mathcal{M}^{(N)*}_{a_{N}\,|\,x_{N}}$ are the parties’ local operations. We then prove that the ICO bound is a measure of the deviation between the map $\mathcal{M}^{*}_{\mathcal{I}}$ and the set of $N$-partite no-signalling channels, that is, the set of quantum processes that do not allow any subset of the parties to transmit information to any other subset (see Methods). Without loss of generality, here we focus our attention on the case where all the coefficients $\alpha_{\vec{a},\vec{x}}$ are nonnegative. In this case, we show that the single-trigger ICO bound is
$\mathcal{I}^{\rm ICO}_{\rm single-trigger}=2^{D_{\max}(\mathcal{M}^{*}_{\mathcal{I}}\,\|\,\mathsf{NoSig})}\,,$ (3)
where $\mathsf{NoSig}$ is the set of $N$-partite no-signalling channels, $D_{\max}(\mathcal{A}\|\mathcal{B}):=\min\{\lambda~{}|~{}\mathcal{A}\leq 2^{\lambda}\,\mathcal{B}\}$ is the max relative entropy between two completely positive maps $\mathcal{A}$ and $\mathcal{B}$ (with the notation $\mathcal{A}\leq\mathcal{B}$ meaning that $\mathcal{B}-\mathcal{A}$ is completely positive), and, for a set of completely positive maps $\mathsf{S}$, $D_{\max}(\mathcal{A}\|\mathsf{S}):=\max_{\mathcal{B}\in\mathsf{S}}D_{\max}(\mathcal{A}\|\mathcal{B})$.
The quantity $D_{\max}(\mathcal{M}^{*}_{\mathcal{I}}\,\|\,\mathsf{NoSig})$ appearing in the ICO bound is also known as the max relative entropy of signalling Chiribella and Ebler (2016). Eq. (3) provides an SDP expression for the ICO bound of all single-trigger causal inequalities. In particular, it yields the ICO bound $\mathcal{I}_{\rm LGYNI}^{\rm ICO}\approx 0.8194$ for the single-trigger causal inequality corresponding to the LGYNI game. This result provides the answer to an open question raised in Ref. Branciard et al. (2015), where the value $\mathcal{I}_{\rm LGYNI}\approx 0.8194$ was obtained by optimizing over local experiments and ICO processes under the assumption that all quantum systems in the parties’ local laboratories are two-dimensional.
Bound for arbitrary causal inequalities. The ICO bound for single-trigger causal inequalities directly implies a general bound for arbitrary causal inequalities. The bound follows from the fact that every correlation $\mathcal{I}$ can be decomposed as a sum of single-trigger correlations, and therefore its ICO bound cannot exceed the sum of the ICO bounds for the single-trigger correlations in its decomposition. Optimizing over all possible decompositions, one obtains the bound
$\mathcal{I}^{\rm ICO}\leq\min_{\left\{\mathcal{I}_{j}\right\}_{j=1}^{n}}\left\{\sum_{j=1}^{n}\,\mathcal{I}_{j}^{\rm ICO}\right\}\,,$ (4)
where $\left\{\mathcal{I}_{j}\right\}_{j=1}^{n}$ are single-trigger correlations satisfying the condition $\sum_{j=1}^{n}\,\mathcal{I}_{j}=\mathcal{I}$.
In Appendix A, we show that the evaluation of the right-hand-side of Eq. (4) is an SDP on a space of fixed dimension, and provide a characterization of this SDP in terms of a set of necessary conditions on quantum correlations with indefinite causal order. The solution of the SDP provides a computable upper bound on the violation of causal inequalities for every desired correlation $\mathcal{I}$.
In the next sections, we show three important implications of our general results.
Maximum quantum violation of the OCB inequality. The first example of a causal inequality is the OCB inequality Oreshkov et al. (2012), which corresponds to a bipartite scenario in which $a_{1},a_{2},$ and $x_{1}$ are bits, while $x_{2}$ is a pair of bits, here denoted by $b$ and $c$, respectively. The inequality arises from a game, in which either Alice is asked to guess Bob’s bit $b$ or Bob is asked to guess Alice’s bit $x_{1}$, depending on the value of bit $c$. When the order between Alice’s and Bob’s experiments is well-defined, only one player can communicate the value of his/her input to the other player, and therefore one of the two players must make a random guess of the other players’ input. Hence, every causal probability distribution satisfies the constraint that the probability that Alice’s guess is correct, plus the probability that Bob’s guess is correct, is upped bounded as ${\mathcal{I}}_{\rm OCB}^{\rm causal}:=P(a_{1}=b\,|\,c=0)+P(a_{2}=x_{1}\,|\,c=1)\leq\frac{3}{2}$, with $P(a_{1}=b\,|\,c=0):=\sum_{x_{1},b}\,p_{A}(b\,|\,x_{1},b,0)/4$ and $P(a_{2}=x_{1}\,|\,c=1):=\sum_{x_{1},b}\,p_{B}(x_{1}\,|\,x_{1},b,1)/4$.
OCB showed that the above causal inequality can be violated by a quantum ICO process and by suitable experiments in Alice’s and Bob’s laboratories, the combination of which reaches the value
${\mathcal{I}}_{\rm OCB}=1+\frac{1}{\sqrt{2}}\,.$ (5)
A fundamental question raised in the original OCB paper is whether higher violations are possible, and, in case they are, what is the maximum violation. The value (5) was shown to be maximum among the violations achievable with a restricted set of local measurements, involving measurement and repreparations with traceless binary observables Brukner (2015). Whether more general types of measurements could lead to higher violations, however, remained as an open question until now. In fact, even the most basic question of whether the violation of the OCB inequality over arbitrary measurements and arbitrary ICO processes can reach its maximum algebraic value $\mathcal{I}_{\rm OCB}=2$ had remained unanswered until the present work.
We now solve the problem in full generality, proving that (5) is indeed the largest violation allowed by arbitrary quantum processes with indefinite causal order and by arbitrary operations in Alice’s and Bob’s laboratory. In fact, we derive the exact ICO bound for a version of the OCB correlation, called the biased OCB correlation Bhattacharya and Banik (2015) and given by
$\mathcal{I}_{{\rm OCB},\alpha}=P(a_{1}=b\,|\,c=0)+\alpha\,P(a_{2}=x_{1}\,|\,c=1)\,,$ (6)
where $\alpha$ is an arbitrary real number. To derive the ICO bound, we observe that the biased OCB correlation is a random mixture of two single-trigger correlations and we evaluate the ICO bound (3) for these two correlations. In this way, we obtain an upper bound on the quantum violation, which we show to coincide with the value found in Bhattacharya and Banik (2015) with a specific example of ICO process (see Appendix B). All together, these results establish the ICO bound
$\mathcal{I}_{\rm OCB,\alpha}^{\rm ICO}=\frac{1+\alpha+\sqrt{1+\alpha^{2}}}{2}\,,$ (7)
thereby providing an analogue of the Tsirelson bound for indefinite causal order.
Figure 1: Geometry of causal, ICO, and general probability distributions. The figure shows the two probabilities appearing in the OCB correlation. The inner square (in green) corresponds to the values achievable by causal probability distributions, while the circle (in blue) corresponds to the values achievable through quantum ICO processes. The outer square (in violet) corresponds to the values achievable by arbitrary, unconstrained probability distributions.
The quantum ICO set of correlations. Our results provide insights into the geometry of the set of probability distributions generated by quantum ICO processes. Let us consider the paradigmatic case of the OCB game, and visualize the possible values of Alice’s and Bob’s success probabilities $P_{A}:=P(a_{1}=b\,|\,c=0)$ and $P_{B}:=P(a_{2}=x_{1}\,|\,c=1)$ by representing them in a two-dimensional plane. With this notation, the ICO bound Eq. (7) is equivalent to the condition $\cos\theta\,(P_{A}-1/2)+\sin\theta\,(P_{B}-1/2)\leq 1/2$ for $\theta=\arctan\alpha$. This condition identifies a circle of radius $1/2$ centred around the point $(1/2,1/2)$, as illustrated in Figure 1. In the figure, we also show the area corresponding to causal probability distributions, which form a square inscribed inside the circle. Finally, arbitrary unconstrained probability distributions occupy the square circumscribing the circle of quantum ICO probabilities.
Intriguingly, Figure 1 coincides with the analogous picture in the case of the CHSH inequality, where the smaller square is the set of probability distributions allowed by local realism, the circle is the set of probability distributions allowed by quantum mechanics, and the larger square is the set of general no-signalling probability distributions Brunner et al. (2014). Later in this paper, we will show that, in the case of causal inequalities, the largest square corresponds to probability distributions achievable in an alternative physical theory where the local parties are restricted to a subset of operations and the processes connecting the parties’ experiments are only required to satisfy the positivity and normalization of probabilities for operations in this restricted subset.
In Appendix C, we provide further insights into the geometry of quantum correlations with indefinite causal order by analyzing the set of correlations arising from the LGYNI game and a biased version thereof.
Bound for Guess Your Neighbor’s Input. Guess Your Neighbor’s input (GYNI) Almeida et al. (2010) is a multipartite quantum game where each party has to guess the input of one of its neighbors. This game is well known in the study of quantum nonlocality, being the first example of a tight Bell inequality with no quantum violation. In the two-party setting, the probability of success is $\mathcal{I}_{\rm GYNI}=P(a_{1}=x_{2},\,a_{2}=x_{1})$. GYNI has also been studied in the context of causal inequalities Branciard et al. (2015); Bavaresco et al. (2019), where it was observed that causal probability distributions satisfy the bound $\mathcal{I}^{\rm causal}_{\rm GYNI}\leq 1/2$ when the settings are uniformly random, while quantum ICO processes can violate this bound. The question about the maximum violation of the GYNI causal inequality has been tackled in Refs. Bavaresco et al. (2019); Kunjwal and Oreshkov (2023). In particular, Ref. Kunjwal and Oreshkov (2023) proved that the quantum ICO violation cannot reach the algebraic maximum $\mathcal{I}_{\rm GYNI}=1$. However, no explicit bound on the maximum violation other than the algebraic maximum has been known up to now.
Using our general expression (4), we can now show that the ICO bound for the GYNI game satisfies the inequality $\mathcal{I}_{\rm GYNI}^{\rm ICO}\leq 0.7592$. At present, this inequality represents the state-of-the-art in upper bounding the quantum ICO violation of the GYNI inequality. Nevertheless, finding the exact value of $\mathcal{I}_{\rm GYNI}^{\rm ICO}$ remains an open problem. Previous numerical results based on a see-saw algorithm using quantum systems of dimension 5 showed that quantum ICO processes can reach the value $\mathcal{I}_{\rm GYNI}\approx 0.6218$ Branciard et al. (2015). An interesting question is whether increasing the dimension of the local quantum systems could yield correlations that go all the way to the upper bound $0.7592$, or whether the maximum quantum violation is attained at a strictly smaller value.
Reaching the algebraic maximum. We have seen that the maximum violation of causal inequalities compatible with standard quantum theory is subject to nontrivial constraints. A natural question is whether it is possible to find alternative physical theories in which the violation can in principle reach its algebraic maximum. In the context of Bell inequalities, this question has been asked by Popescu and Rorlich Popescu and Rohrlich (1994), who showed that the algebraic maximum of the CHSH inequality can be achieved in an alternative theory, now known as boxworld Gross et al. (2010). For causal inequalities, we provide two examples of theories that reach the algebraic maximum of the GNYI inequality and other two-party causal inequalities.
Consider a variant of quantum theory in which the experimenters are restricted to perform only bistochastic instruments, that is, instruments $(\mathcal{M}_{a\,|\,x})_{a=1}^{m}$ satisfying the condition that the map $\sum_{a}\mathcal{M}_{a\,|\,x}$ is both trace-preserving and identity preserving. We call this theory bistochastic quantum theory Chiribella et al. (2021). As in ordinary quantum theory, one can define the set of all logically conceivable supermaps in bistochastic quantum theory. These supermaps, introduced in Ref. Chiribella and Liu (2022), represent the most general processes that could in principle connect the local operations performed by parties obeying to bistochastic quantum theory, giving rise to probability distributions as in Eq. (2), where now all the instruments are required to be bistochastic. In the two-party scenario, an explicit characterization of the admissible supermaps is provided in Methods.
The above construction also applies to the bistochastic version of classical theory, which can be obtained from bistochastic quantum theory by subjecting every system to a decoherence map. Explicitly, a bistochastic classical channel is described by a conditional probability distribution specifying the probability $q^{(i)}(s_{i}^{\prime}\,|\,s_{i})\geq 0$ that an input state $s_{i}$ at time $t_{i}$ is transformed into an output state $s_{i}^{\prime}$ at time $t_{i}^{\prime}$, and satisfying the conditions $\sum_{s_{i}^{\prime}}q^{(i)}(s_{i}^{\prime}\,|\,s_{i})=1$ and $\sum_{s_{i}}q^{(i)}(s_{i}^{\prime}\,|\,s_{i})=1$. In classical bistochastic theory, the most general experiments performed by the $i$-th party correspond to classical bistochastic instruments, mathematically described by subnormalized probability distributions $q^{(i)}_{a_{i}\,|\,x_{i}}(s_{i}^{\prime}\,|\,s_{i})$ satisfying the condition that $\sum_{a_{i}}q^{(i)}_{a_{i}\,|\,x_{i}}(s_{i}^{\prime}\,|\,s_{i})$ is a classical bistochastic channel for every setting $x_{i}$. The set of logically conceivable supermaps in this theory coincides with the decohered version of the supermaps allowed in bistochastic quantum theory (see Methods and Appendix D for more details).
In the following, the maximum of a correlation $\cal I$ over the dimensions of the parties’ systems, over all choices of bistochastic instruments, and over all admissible supermaps will be called the bistochastic ICO (BICO) value. The main result of this section is that the quantum and classical BICO values coincide and are equal to the algebraic maximum for every two-party correlation with up to three settings. To prove this result, we start from the GYNI game, which can be won with certainty only if Alice and Bob can perfectly signal to one another.
Consider a scenario where all the systems entering and exiting the parties’ laboratories are classical bits. To win the GYNI game, Alice and Bob adopt the following strategy: First, they measure the values of the bits $s_{1}$ and $s_{2}$ entering their laboratories, respectively, and, depending on their settings, they either leave the bit unchanged or they filp them, obtaining the new values $s_{1}^{\prime}=s_{1}\oplus x_{1}$ and $s_{2}^{\prime}=s_{2}\oplus x_{2}$, respectively (here $\oplus$ denotes addition modulo 2). The final value of the bit then becomes the outcome of their measurements, namely $a_{1}=s_{1}^{\prime}$ and $a_{2}=s_{2}^{\prime}$. Mathematically, this strategy corresponds to the classical bistochastic instruments $q^{(1)}_{a_{1}\,|\,x_{1}}(s_{1}^{\prime}\,|\,s_{1})=\delta_{s_{1}^{\prime},s_{1}\oplus x_{1}}\,\delta_{a_{1},s_{1}^{\prime}}$ for Alice and $q^{(2)}_{a_{2}\,|\,x_{2}}(s_{2}^{\prime}\,|\,s_{2})=\delta_{s_{2}^{\prime},s_{2}\oplus x_{2}}\,\delta_{a_{2},s_{2}^{\prime}}$ for Bob. To achieve unit probability of winning the GYNI game, Alice’s and Bob’s local experiments are connected by a deterministic process $\mathcal{S}$, represented as a function that maps the final values of the bits $(s_{1}^{\prime},s_{2}^{\prime})$ into their initial values $(s_{1},s_{2})$ according to the rule
$s_{1}=s_{1}^{\prime}\oplus s_{2}^{\prime}\qquad{\rm and}\qquad s_{2}=s_{1}^{\prime}\oplus s_{2}^{\prime}\,.$ (8)
The action of the process $\mathcal{S}$ on the parties’ operations produces the probability distribution $\mathcal{S}(q^{(1)}_{a_{1}\,|\,x_{1}},q^{(2)}_{a_{2}\,|\,x_{2}})$ $=\sum_{s_{1}^{\prime},s_{2}^{\prime}}\,q^{(1)}_{a_{1}\,|\,x_{1}}(s_{1}^{\prime}\,|\,s_{1}^{\prime}\oplus s_{2}^{\prime})\,q^{(2)}_{a_{2}\,|\,x_{2}}(s_{2}^{\prime}\,|\,s_{1}^{\prime}\oplus s_{2}^{\prime})$. In Methods, we show that the process $\mathcal{S}$ is admissible, in the sense that it gives rise to a normalized probability distribution for every choice of bistochastic instruments $q^{(1)}_{a_{1}\,|\,x_{1}}$ and $q^{(2)}_{a_{2}\,|\,x_{2}}$. For the specific choice of instruments described earlier in this paragraph, the process $\mathcal{S}$ generates the perfect signalling distribution $p(a_{1},a_{2}\,|\,x_{1},x_{2})=\delta_{a_{1},x_{2}}\,\delta_{a_{2},x_{1}}$, which allows Alice and Bob to win the GYNI game with certainty.
In Appendix E, we extend the above result from GYNI to arbitrary causal inequalities with two parties and up to three settings. Specifically, we prove that perfect two-way signalling of ternary digits is achievable by choices of classical bistochastic instruments and classical supermaps. Equipped with a perfect two-way signalling of each other’s setting, the two parties can generate arbitrary deterministic conditional probability distributions, and therefore can reach the algebraic maximum for every given correlation function. In addition, we show that if each party has only two settings, then every bipartite conditional probability distribution can be generated by a suitable choices of instruments and supermaps in classical bistochastic theory. Since instruments and processes in bistochastic classical theory are special cases of instruments and processes in bistochastic quantum theory, the same results hold in the quantum case.
III Discussion
Our general bound on the violation of causal inequalities is based on an SDP relaxation of the original problem of computing the quantum ICO bound. For certain correlations, such as the OCB correlation and all single-trigger correlations, we have shown that the SDP gives the exact value of the quantum ICO bound. In principle, one could ask whether our bound is tight for all possible causal inequalities. The analogy with Bell inequalities, however, suggests a negative answer. In Bell scenarios, a converging sequence of upper bounds on the value of maximal quantum violations is provided by the Navascués-Pironio-Acín SDP hierarchy Navascués et al. (2007, 2008). The analogy with this situation suggests that our SDP relaxation may be just the first level of a a similar hierarchy of SDPs. Determining whether this analogy is correct, and, in the affirmative case, identifying the other levels of the hierarchy are among the most important research directions opened by our work. Another interesting direction is to extend our method for the calculation of the ICO bound to other type of inequalities with non-trivial causal structure, such as the inequalities recently studied in Refs. Gogioso and Pinzani (2023); van der Lugt et al. (2023).
A natural development of our research is to establish self-testing results for causal inequalities, in analogy to self-testing for Bell inequalities Mayers and Yao (1998, 2004). For example, it is interesting to determine whether the OCB process is the only quantum process (up to local transformations) that achieves the maximum violation of the OCB inequality. More broadly, it is interesting to search for applications of the violation of causal inequalities to quantum technologies, in analogy to the applications of the violation of Bell inequalities in in quantum cryptography Acin et al. (2006); Colbeck (2009); Pironio et al. (2010), communication complexity Buhrman et al. (2010), and control of quantum hardware Reichardt et al. (2013). While the physical realization of general quantum ICO processes is still an open problem, exploring the potential applications of causal inequalities contributes to a deeper information-theoretic understanding of indefinite causal order in quantum theory.
Finally, our results open up a search for physical principles capable of explaining why the violation of causal inequalities by quantum ICO processes is not equal, in general, to the algebraic maximum, and, of determining the exact value of the quantum violation. In the context of Bell inequalities, the analogue question was originally raised by Popescu and Rohrlich Popescu and Rohrlich (1994), and led to the discovery of new information theoretic principles, such as non-trivial communication complexity Van Dam (1999); Brassard et al. (2006); Brunner and Skrzypczyk (2009), non-trivial nonlocal computation Linden et al. (2007), information causality Pawłowski et al. (2009), macroscopic locality Navascués and Wunderlich (2010), and local orthogonality Fritz et al. (2013). The formulation of new principles that capture the quantum violation of causal inequalities is likely to be valuable for the understanding of quantum physics in spacetime, and may eventually lead to a new formulation of quantum theory that does not require a pre-defined causal structure. In turn, such a formulation may offer guidance in the long-standing problem of unifying quantum theory with Einstein’s theory of relativity.
IV Methods
Labelled projective instruments. We now develop a way to reduce the search over the set of all possible local experiments to a search over a smaller set, generated by ideal measurements and ideal state preparations. This reduction plays a similar role as the reduction to projective measurements in Bell scenarios.
In the standard framework of quantum theory, an experiment with $m$ possible outcomes is described by a quantum instrument Davies and Lewis (1970); Ozawa (1984); Heinosaari and Ziman (2011), that is, a collection of completely positive maps $(\mathcal{M}_{a})_{a=0}^{m-1}$ satisfying the condition that their sum is trace-preserving. Each map $\mathcal{M}_{a}$ describes a physical process transforming an input system with Hilbert space $\mathcal{H}_{\rm in}$ into a (possibly different) output system, with Hilbert space $\mathcal{H}_{\rm out}$. In the special case $\mathcal{H}_{\rm in}=\mathcal{H}_{\rm out}$, an instrument $(\mathcal{M}_{a})_{a=0}^{n-1}$ is called projective if each map $\mathcal{M}_{a}$ is of the form $\mathcal{M}_{a}(\rho)=P_{a}\rho P_{a}$, where $P_{a}$ is a projector and the projectors $(P_{a})_{a=0}^{m-1}$ form a resolution of the identity.
We now introduce the notion of labelled projective instruments, that is, instruments that output a label describing their settings. For every setting $x$, let $(\mathcal{N}_{a\,|\,x})_{a=0}^{m-1}$ be a projective instrument on a quantum system with Hilbert space $\mathcal{H}_{\rm in}$, and let $(\rho_{x})_{x=0}^{n-1}$ be a set of perfectly distinguishable density matrices for an auxiliary quantum system with Hilbert space $\mathcal{H}_{\rm aux}$. A labelled projective instrument is an instrument $(\mathcal{M}_{a\,|\,x})_{a=0}^{m-1}$ of the form $\mathcal{M}_{a\,|\,x}=\mathcal{N}_{a\,|\,x}\otimes\rho_{x}$.
Our first result is that the optimization of a conditional probability distribution $p(\vec{a}\,|\,\vec{x})$ over arbitrary instruments and arbitrary ICO processes can be restricted without loss of generality to an optimization over labelled projective instruments. The key result is the following theorem:
Theorem 1
For every set of quantum instruments $\big{(}\mathcal{M}^{(i)}_{a_{i}|x_{i}}\big{)}_{a_{i}=0}^{m_{i}-1}$, $x_{i}\in\{0,\dots,n_{i}-1\}$, $i\in\{1,\dots,N\}$ and every process $\mathcal{S}$ acting on them, there exists a set of labelled projective instruments $\big{(}\mathcal{M}^{(i)\prime}_{a_{i}|x_{i}}\big{)}_{a_{i}=0}^{m_{i}-1}$, $x_{i}\in\{0,\dots,n_{i}-1\}$, $i\in\{1,\dots,N\}$ (with possibly different input and output systems) and a process $\mathcal{S}^{\prime}$ acting on them such that
$\mathcal{S}(\mathcal{M}^{(1)}_{a_{1}|x_{1}},\dots,\mathcal{M}^{(N)}_{a_{N}|x_{N}})=\mathcal{S}^{\prime}(\mathcal{M}^{(1)\prime}_{a_{1}|x_{1}},\dots,\mathcal{M}^{(N)\prime}_{a_{N}|x_{N}})$ (9)
for all settings $\vec{x}$ and all outcomes $\vec{a}$. In addition, the labelled projective measurements associated to each party $i$ can be chosen without loss of generality to have projectors of the same rank $r_{i}$ for every outcome $a_{i}$ and every setting $x_{i}$.
The proof uses Ozawa’s dilation theorem Ozawa (1984) for quantum instruments, combined with a use of auxiliary systems to include unitary state changes depending on the settings. The details of the proof are provided in Appendix F.
Theorem 1 can be used in every problem involving the joint optimization of local instruments and global ICO processes. In particular, it can be used to simplify the search for the ICO bound of arbitrary causal inequalities. In the next section, we will show that a strengthening of Theorem 1 can be provided in the case of single-trigger causal inequalities.
The canonical instrument. We now show that the ICO bound for single-trigger correlations can be achieved by the canonical choice of labelled projective instruments $\mathcal{M}_{a_{i}\,|\,x_{i}}^{(i)*}$ provided in the main text.
In general, the calculation of the ICO bound involves a double optimization, over the all possible local instruments and over all possible ICO processes. As an intermediate step, it is useful to consider the simpler problem where the local instruments are fixed and the optimization runs over the set of ICO processes. Using Eqs. (1) and (2), the correlation achieved by an ICO process $\mathcal{S}$ can be written as $\mathcal{I}=\mathcal{S}(\mathcal{M}_{\mathcal{I}})$, with $\mathcal{M}_{\mathcal{I}}=\sum_{\vec{a},\vec{x}}\alpha_{\vec{a},\vec{x}}\,\mathcal{M}_{a_{1}\,|\,x_{1}}^{(1)}\otimes\cdots\otimes\mathcal{M}_{a_{N}\,|\,x_{N}}^{(N)}$. The maximum correlation achieved by arbitrary processes is given by $\upsilon(\mathcal{M}_{\mathcal{I}}):=\max_{\mathcal{S}}\mathcal{S}(\mathcal{M}_{\mathcal{I}})$. The maximization is a semidefinite program, whose solution can be equivalently computed as Chiribella and Ebler (2016)
$\upsilon(\mathcal{M}_{\mathcal{I}})=\min\Big{\{}\eta\in\mathbb{R}~{}|~{}\exists\mathcal{C}\in\operatorname{Aff}(\mathsf{NoSig}):\,\eta\,\mathcal{C}\geq\mathcal{M}_{\mathcal{I}}\Big{\}}\,,$ (10)
where $\operatorname{Aff}(\mathsf{NoSig})$ is the set of affine combinations of no-signalling channels, and, for two linear maps $\mathcal{A}$ and $\mathcal{B}$ with the same input and output spaces, $\mathcal{A}\geq\mathcal{B}$ means that $\mathcal{A}-\mathcal{B}$ is completely positive. When the coefficients $\alpha_{\vec{a},\vec{x}}$ are nonnegative, the minimization can be restricted without loss of generality to the set of no-signalling channels, and the minimum is given $2^{D_{\max}(\mathcal{M}_{\mathcal{I}}\|\mathsf{NoSig})}$, where $D_{\max}(\mathcal{M}_{\mathcal{I}}\|\mathsf{NoSig})$ is the max relative entropy distance between the operator $M_{\mathcal{I}}$ and the set of no-signalling channels. In short, the maximum correlation achievable with a fixed set of local instruments is given by the deviation of the map $\mathcal{M}_{\mathcal{I}}$ from the set of no-signalling channels.
We are now ready to tackle the full problem of computing the ICO value. With the above notation, the ICO value can be written as $\mathcal{I}^{\rm ICO}=\max_{\mathcal{M}_{\mathcal{I}}}\upsilon(\mathcal{M}_{\mathcal{I}})$, where the maximization is over all maps $\mathcal{M}_{\mathcal{I}}$ generated by all possible local quantum instruments with inputs and outputs of arbitrary dimensions. By Theorem 1, the maximization can be restricted without loss of generality to maps $\mathcal{M}_{\mathcal{I}}$ generated by labelled projective instruments. In Appendix G, we show that, for single-trigger correlations and labelled projective instruments, the map $\mathcal{M}_{\mathcal{I}}$ can be decomposed into a convex combination of maps associated to instruments of an even simpler form, as stated by the following theorem:
Theorem 2
Every map $\mathcal{M}_{\mathcal{I}}$ associated to a given single trigger correlation and a given set of labelled projective instruments can be decomposed into a convex combination $\mathcal{M}_{\mathcal{I}}=\sum_{j}\,p_{j}\,\mathcal{M}_{\mathcal{I},j}$, where $(p_{j})$ is a probability distribution and, for every $j$, $\mathcal{M}_{\mathcal{I},j}$ is the map associated to local instruments of the form
$\mathcal{M}_{\,a_{i}|x_{i}}^{(i,j)}(\rho)=\begin{cases}P_{a_{i},\xi_{i}}^{(i)}\,\rho\,P_{a_{i},\xi_{i}}^{(i)}\otimes|\xi_{i}\rangle\langle\xi_{i}|&x_{i}=\xi_{i}\,,\\ \frac{1}{m_{i}}\,U_{x_{i}}^{(i,j)}\,\rho\,U_{x_{i}}^{(i,j){\dagger}}\otimes|x_{i}\rangle\langle x_{i}|&x_{i}\neq\xi_{i}\,.\end{cases}$ (11)
where $\big{(}P_{a_{i}|x_{i}}^{(i)}\big{)}_{a_{i}=0}^{m_{i}-1}$ are the projectors appearing in the original labelled projective instrument, and $U_{x_{i}}^{(i,j)}$ is a unitary operator for every $i,j,$ and $x_{i}$.
Operationally, the instruments appearing in Theorem 2 describe experiments where the $i$-th player performs a labelled projective instrument when their setting is the trigger, and, for all the other settings, the player generates a uniformly random outcome, while performing a unitary gate on the system entering their laboratory. Since $\upsilon$ is a convex function of the map $\mathcal{M}_{\mathcal{I}}$, Theorem 2 guarantees that its maximization can be restricted without loss of generality to maps generated by instruments of this form. Note that the canonical instruments $(\mathcal{M}^{(i)*})_{a_{i}|x_{i}}$ are a special case of instrument of the form (11), corresponding to the case where all projectors are rank-one, and all unitary operators are equal to the identity operator.
In Appendix H we show that every value of the function $\upsilon(\mathcal{M}_{\mathcal{I}})$ achievable by instruments of the form (11) can be achieved by the canonical instrument. Intuitively, the argument is as follows: first, Theorem 1 guarantees that all the instruments associated to the same party have projectors of the same rank. For single-trigger correlations, Theorem 2 ensures that one can effectively consider only one projective instrument per party. But then, a single projective instrument with projectors of equal rank can be reduced to a rank-one projective instrument by appending an additional quantum system in the party’s local laboratory, without performing any operation on it. Without loss of generality, the additional system can be incorporated in the definition of the process $\mathcal{S}$ connecting the local laboratories. Similarly, arbitrary unitary operations depending on the parties’ settings can be obtained from the identity operation, by appending to the process $\mathcal{S}$ a controlled unitary operation, controlled by the state of the auxiliary systems that carry the values of the parties’ settings (see Appendix H for the details). The above argument implies the bound $\upsilon(\mathcal{M}_{\mathcal{I}}^{*})\geq\upsilon(\mathcal{M}_{\mathcal{I}})$ for every map $\mathcal{M}_{\mathcal{I}}$ generated by local instruments, and therefore $\mathcal{I}^{\rm ICO}_{\rm single-trigger}=\upsilon(\mathcal{M}_{\mathcal{I}}^{*}).$
Using Eq. (10) we then obtain an explicit SDP expression for the ICO value. When the coefficients $\alpha_{\vec{a},\vec{x}}$ are nonnegative, this expression reduces to the max relative entropy distance, in agreement with Eq. (3) in the main text.
Admissible supermaps in bistochastic quantum and classical theory. Here we consider time-symmetric versions of quantum and classical theories, in which local experiments are described by bistochastic instruments, and we characterize the set of all logically conceivable supermaps. In the quantum case, this set of supermaps was introduced in Ref. Chiribella and Liu (2022) and were shown to be in one-to-one correspondence with a subset of positive semidefinite operators. In the two-party case, the logically conceivable supermaps admit a simple characterization in terms of their Choi operators: denoting by $A_{\rm in}$ and $A_{\rm out}$ ($B_{\rm in}$ and $B_{\rm out}$) the input and output systems of the first (second) party, an admissible supermap is described by an operator $S\in L(\mathcal{H}_{A_{\rm in}}\otimes\mathcal{H}_{B_{\rm out}}\otimes\mathcal{H}_{B_{\rm in}}\otimes\mathcal{H}_{B_{\rm out}})$ satisfying the conditions (see Appendix D for a proof)
$S\geq 0\,,$
$\operatorname{Tr}(S)=d_{A}d_{B}\,,$
${}_{[(1-A_{\rm in})(1-A_{\rm out})B_{\rm in}B_{\rm out}]}S=0\,,$ (12)
${}_{[(1-B_{\rm in})(1-B_{\rm out})A_{\rm in}A_{\rm out}]}S=0\,,$
${}_{[(1-A_{\rm in})(1-A_{\rm out})(1-B_{\rm in})(1-B_{\rm out})]}S=0\,,$
where $\mathcal{H}_{X}$ denotes the Hilbert space of system $X$, and $L(\mathcal{H})$ represents the set of linear operators on the Hilbert space $\mathcal{H}$. We use the notation ${}_{[X]}S:=\operatorname{Tr}_{X}[S]\otimes\frac{I_{X}}{d_{X}}$ for a system $X$ of dimension $d_{X}$, and the shorthand ${}_{[\sum\alpha_{X}X]}S:=\sum_{X}\alpha_{X}\cdot{}_{[X]}S$ for a linear combination of the terms $\{{}_{[X]}S\}_{X}$ associated with a collection of systems indexed by $X$. For instance, ${}_{[(1-X)Y]}S$ serves as a shorthand for ${}_{[Y]}S-{}_{[XY]}S$. When the parties perform local instruments, respectively, the probability in Eq. (2) is given by $p(a_{1},a_{2}|x_{1},x_{2})=\operatorname{Tr}\left[S\,(M^{(1)}_{a_{1}|x_{1}}\otimes M^{(2)}_{a_{2}|x_{2}})\right]$.
Admissible supermaps in bistochastic classical theories can be obtained from their quantum counterparts by fixing an orthonormal basis for every local Hilbert space and setting all the off-diagonal elements of the Choi operator $S$ to zero. In the bipartite case, an admissible supermap is described by a matrix with entries $\sigma_{\mathbf{s}\mathbf{s}^{\prime}}=\langle s_{1}|\langle s_{1}^{\prime}|\langle s_{2}|s_{2}^{\prime}|S|s_{1}\rangle|s_{1}^{\prime}\rangle|s_{2}\rangle|s_{2}^{\prime}\rangle$ where, for $j\in\{1,2\}$, $\{|s_{j}\rangle\}$ and $\{|s_{j}^{\prime}\rangle\}$ are bases for the Hilbert spaces of the input and output spaces of the $j$-th party. In particular, the supermap described in the main text corresponds to the matrix with entries $\sigma_{\mathbf{s}\mathbf{s}^{\prime}}=\delta_{s_{1},s_{1}^{\prime}\oplus s_{2}^{\prime}}\delta_{s_{2},s_{1}^{\prime}\oplus s_{2}^{\prime}}$, which satisfies all the constraints in Eq. (12) and therefore represents a logically admissible process in bistochastic quantum and classical theory.
Data availability
The authors declare that the data supporting the findings of this study are available within the paper and in the supplementary information files.
Acknowledgments
The authors thank R. Ramanathan, Č. Brukner, M. Araujo, O. Oreshokov, R. Kunjwal and S. Yoshida for helpful comments. This work has been supported by the Hong Kong Research Grant Council through the Senior Research Fellowship Scheme SRFS2021-7S02 and through the Research Impact Grant R7035-21F, and by the F.R.S.-FNRS under project CHEQS within the Excellence of Science (EOS) program. The work was also supported by the John Templeton Foundation through grant 62312, The Quantum Information Structure of Spacetime (qiss.fr). The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. Research at the Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Research, Innovation and Science.
Author contributions
ZL introduced the notion of single-trigger correlations, derived the bounds on quantum ICO correlations, and showed the achievability of the algebraic maximum by processes in bistochastic classical and quantum theories. GC supervised the research and proved the reduction to labelled projective measurements (Theorem 1). GC and ZL wrote the manuscript.
Competing interests
The authors declare no competing interests.
Appendix A Semidefinite programming (SDP) relaxation of ICO bounds
In this section, we introduce Choi-Jamiołkowski isomorphism which allows us to represent quantum processes and local quantum operations as operators; then we explicitly present the SDPs for computing the ICO bound of single-trigger correlations and for computing upper bounds of general correlations.
Choi-Jamiołkowski isomorphism
Through Choi-Jamiołkowski isomorphism Jamiołkowski (1972); Choi (1975), a linear map $\mathcal{K}$ on operators can be conveniently represented as a Choi operator, defined as follows:
$\operatorname{Choi}(\mathcal{K}):=\sum_{i,j}|i\rangle\langle j|\otimes\mathcal{K}(|i\rangle\langle j|)\,.$ (13)
If $\mathcal{K}$ is a completely positive map with Kraus operators $\{K_{i}\}$, it holds that
$\operatorname{Choi}(\mathcal{K})=\sum_{i}|K_{i}\rangle\!\rangle\langle\!\langle K_{i}|\,,$ (14)
where we use the double-ket notation defined as $|M\rangle\!\rangle=\sum_{j}|j\rangle\otimes M|j\rangle$.
Quantum supermaps and process matrices
In an experiment without assumption of global causal structure, the local operations are described by quantum instruments, that is, collections of linear maps $(\mathcal{M}_{j})_{j}$ labelled by possible measurement outcomes $j$. Each linear map $\mathcal{M}_{j}$, called a quantum operation, transforms density matrices on an input system into (generally subnormalized) density matrices on an output system, and is required to be completely positive. In the following, we will denote the $i$-th party’s instruments (conditional on the setting $x_{i}$) by $\left(\mathcal{M}^{(i)}_{a_{i}\mid x_{i}}\right)_{a_{i}}$, each map $\mathcal{M}^{(i)}_{a_{i}\mid x_{i}}$ transforming quantum states of an input system $A_{\rm I}^{(i)}$ into quantum states of an output system $A_{\rm O}^{(i)}$. The most general rule for assigning probabilities to outcomes is provided by a process $\mathcal{S}$ that transforms all local quantum operations $(\mathcal{M}^{(1)},\cdots,\mathcal{M}^{(N)})$ into probabilities $\mathcal{S}(\mathcal{M}^{(1)},\cdots,\mathcal{M}^{(N)})$. In this way, the joint conditional probability distribution of outcomes is given by
$p(\vec{a}\mid\vec{x})=\mathcal{S}\left(\mathcal{M}^{(1)}_{a_{1}\mid x_{1}},\cdots,\mathcal{M}^{(N)}_{a_{N}\mid x_{N}}\right)\,.$ (15)
Mathematically, a process is a completely positive (CP) map transforming a list of channels to the unit probability, which is a special case of quantum supermaps, the physically admissible operations transforming quantum channels Chiribella et al. (2008, 2009b). Given a supermap $\mathcal{S}$, we can induce a map $\widehat{\mathcal{S}}$ which transforms the Choi operators of channels,
$\widehat{\mathcal{S}}:\quad\operatorname{Choi}(\mathcal{C})\mapsto\operatorname{Choi}(\mathcal{S}(\mathcal{C}))\,,$ (16)
and we define the Choi operator of $\mathcal{S}$ to be the Choi operator of $\widehat{\mathcal{S}}$, i.e. $\operatorname{Choi}(\mathcal{S}):=\operatorname{Choi}(\widehat{\mathcal{S}})$. The condition that $\mathcal{S}$ is a CP map is equivalent to the condition that $\operatorname{Choi}(\mathcal{S})$ is a postive operator, i.e. $\operatorname{Choi}(\mathcal{S})\geq 0$.
The constraints on a process $\mathcal{S}$ that are in principle compatible with the validity of quantum mechanics in all local laboratories can be completely characterized in terms of the Choi operator $S$ of the process $\mathcal{S}$, called the process matrix Oreshkov et al. (2012), satisfying the conditions (i) $S\geq 0$; (ii) $S\in\operatorname{DualAff}(\operatorname{Choi}(\mathsf{NoSig}))$ where $\operatorname{Choi}(\mathsf{NoSig}))$ is the set of Choi operators of no-signalling channels and $\operatorname{DualAff}$ denotes the dual affine space. See Refs. Oreshkov et al. (2012); Araújo et al. (2015) for an explicit characterization of process matrices. The joint conditional probability in Eq. (15) is then computed by
$p(\vec{a}\mid\vec{x})=\operatorname{Tr}\left[S^{T}\cdot\bigotimes_{i=1}^{N}M_{a_{i}\mid x_{i}}^{(i)}\right]\,.$ (17)
where $M_{a_{i}\mid x_{i}}^{(i)}$ are Choi operators of the corresponding instruments $\mathcal{M}_{a_{i}\mid x_{i}}^{(i)}$ and $S^{T}$ represents transposition of the process matrix in the computational basis.
SDP duality
In Methods, with fixed instruments $\left(\mathcal{M}^{(i)}_{a_{i}\mid x_{i}}\right)_{a_{i}}$, we associate every correlation $\mathcal{I}=\sum_{\vec{a},\vec{x}}\alpha_{\vec{a},\vec{x}}\,p(\vec{a}\mid\vec{x})$ with a map $\mathcal{M}_{\mathcal{I}}$. The Choi operator of $\mathcal{M}_{\mathcal{I}}$ will be referred to as the performance operator, which is given by
$\Omega_{\mathcal{I}}:=\operatorname{Choi}(\mathcal{M}_{\mathcal{I}})=\sum_{\vec{a},\vec{x}}\alpha_{\vec{a},\vec{x}}\,\bigotimes_{i=1}^{N}\operatorname{Choi}(\mathcal{M}_{a_{i}\mid x_{i}}^{(i)})\,,$ (18)
In particular, when $\mathcal{I}$ is a single-trigger correlation with triggers $\vec{\xi}$, we associated it with a single-trigger operator, which is the performance operator $\Omega_{\mathcal{I}}^{*}$ with the canonical choice of instruments, and is in turn the Choi operator of the map $\mathcal{M}_{\mathcal{I}}^{*}$ in the main text. The single-trigger operator $\Omega_{\mathcal{I}}^{*}$ is given by
$\Omega_{\mathcal{I}}^{*}=\sum_{\vec{a},\vec{x}}\alpha_{\vec{a},\vec{x}}\,P_{\vec{\xi},\vec{a},\vec{x}}$ (19)
where $P_{\vec{\xi},\vec{a},\vec{x}}$ are projectors defined as
$P_{\vec{\xi},\vec{a},\vec{x}}:=\bigotimes_{i=1}^{N}P_{\xi_{i},a_{i},x_{i}}^{(i)}$ (20)
with
$P_{\xi_{i},a_{i},x_{i}}^{(i)}=\begin{cases}|a_{i}\rangle\langle a_{i}|\otimes|a_{i}\rangle\langle a_{i}|\otimes|\xi_{i}\rangle\langle\xi_{i}|&x_{i}=\xi_{i}\,,\\ |\Phi\rangle\langle\Phi|\otimes|x_{i}\rangle\langle x_{i}|&x_{i}\neq\xi_{i}\,,\end{cases}$ (21)
where $|\Phi\rangle:=\sum_{j}|j\rangle\otimes|j\rangle/\sqrt{m_{i}}$ is the canonical maximally entangled state.
With respective to a performance operator $\Omega_{\mathcal{I}}$, the optimization of the corresponding correlation $\mathcal{I}$ can be expressed through the following SDP Chiribella and Ebler (2016):
maximize $\operatorname{Tr}(\Omega_{\mathcal{I}}S)$ (22)
subject to $S\in\operatorname{DualAff}(\operatorname{Choi}(\mathsf{NoSig}))$
$S\geq 0\,.$
The dual problem of (22) is
minimize $\eta$ (23)
subject to $\eta C\geq\Omega_{\mathcal{I}}$
$C\in\operatorname{Aff}(\operatorname{Choi}(\mathsf{NoSig}))\,,$
where $\operatorname{Aff}(\operatorname{Choi}(\mathsf{NoSig}))$ denotes the set of affine combinations of the Choi operators of no-signalling channels.
By the property of duality, every dual feasible point $(\eta,C)$ provides an upper bound $\eta$ of the SDP pair (22, 23). Furthermore, strong duality holds because both primal optimal value and dual optimal value are finite and the primal feasible set contains a strictly positive operator (the process matrix proportional to the identity) Watrous (2018); Chiribella and Ebler (2016). In particular, the dual problem (23), with the performance operator being a single-trigger operator $\Omega_{\mathcal{I}}^{*}$, is the SDP corresponding to the ICO bound of the corresponding single-trigger correlation $\mathcal{I}$. When the coefficients of a single-trigger correlation $\mathcal{I}$ are non-negative, the corresponding single-trigger operator $\Omega_{\mathcal{I}}^{*}$ is positive and thus the dual problem (23) is equivalent to $\min_{C\in\operatorname{Choi}(\mathsf{NoSig})}\{\eta\in\mathbb{R}\mid\Omega_{\mathcal{I}}^{*}\leq\eta C\}=2^{D_{\max}(\Omega_{\mathcal{I}}^{*}\|\operatorname{Choi}(\mathsf{NoSig}))}$, which is in turn equivalent to the formula $\mathcal{I}^{\rm ICO}=2^{D_{\max}(\mathcal{M}_{\mathcal{I}}^{*}\|\mathsf{NoSig})}$ in the main text.
SDP relaxation in an explicit form
Now we present an SDP relaxation of the problem of computing the ICO bound in an explicit form. For given triggers $\vec{\xi}$, define the following operator for every setting vector $\vec{x}$ and outcome vector $\vec{a}$:
$Q_{\vec{\xi},\vec{a},\vec{x}}:=\bigotimes_{i=1}^{N}Q_{\xi_{i},a_{i},x_{i}}^{(i)}$ (24)
with
$Q_{\xi_{i},a_{i},x_{i}}^{(i)}=\begin{cases}|a_{i}\rangle\langle a_{i}|\otimes|a_{i}\rangle\langle a_{i}|\otimes|\xi_{i}\rangle\langle\xi_{i}|&x_{i}=\xi_{i}\,,\\ \frac{1}{m_{i}}|\Phi\rangle\langle\Phi|\otimes|x_{i}\rangle\langle x_{i}|&x_{i}\neq\xi_{i}\,.\end{cases}$ (25)
By the definition of single-trigger operator, a single-trigger operator $\Omega_{\vec{\xi}}$ with triggers $\vec{\xi}$ is a linear combination of the projectors $\left\{P_{\vec{\xi},\vec{a},\vec{x}}\right\}_{\vec{a},\vec{x}}$, which is in turn equivalent to the following constraint:
$\Omega_{\vec{\xi}}=\sum_{\vec{a},\vec{x}}\operatorname{Tr}[\Omega_{\vec{\xi}}~{}Q_{\vec{\xi},\vec{a},\vec{x}}]P_{\vec{\xi},\vec{a},\vec{x}}\,,$ (26)
and the coefficients of the corresponding single-trigger correlation are given by
$\alpha^{\vec{\xi}}_{\vec{a},\vec{x}}=\operatorname{Tr}[\Omega_{\vec{\xi}}~{}Q_{\vec{\xi},\vec{a},\vec{x}}]\,.$ (27)
Combining Eq. (26), Eq. (27) and the dual SDP (23) for single-trigger correlations, the optimization over the decomposition of a general correlation $\mathcal{I}=\sum_{\vec{a},\vec{x}}\alpha_{\vec{a},\vec{x}}\,p(\vec{a}\mid\vec{x})$ into single-trigger correlations can be formulated into the following SDP:
minimize $\frac{\sum_{\vec{\xi}}\operatorname{Tr}(C_{\vec{\xi}})}{\prod_{i=1}^{N}m_{i}}$ (28)
subject to $\forall\vec{\xi}\quad\Omega_{\vec{\xi}}=\sum_{\vec{a},\vec{x}}\operatorname{Tr}[\Omega_{\vec{\xi}}~{}Q_{\vec{\xi},\vec{a},\vec{x}}]P_{\vec{\xi},\vec{a},\vec{x}}\,,$
$\forall\vec{a}~{}\forall\vec{x}\quad\alpha_{\vec{a},\vec{x}}=\sum_{\vec{\xi}}\operatorname{Tr}[\Omega_{\vec{\xi}}~{}Q_{\vec{\xi},\vec{a},\vec{x}}]\,,$
$\forall\vec{\xi}\quad C_{\vec{\xi}}\in\operatorname{Span}(\operatorname{Choi}(\mathsf{NoSig}))\,,$
$\forall\vec{\xi}\quad C_{\vec{\xi}}\geq\Omega_{\vec{\xi}}\,.$
Necessary conditions on the quantum ICO set
From another point of view, the ICO bounds of single-trigger correlations are equivalent to a set of necessary conditions on the quantum ICO set. To characterize these necessary conditions, we introduce a projection operator, denoted as $\Pi_{\vec{\xi}}$, which acts on the space of conditional probability distributions. This projection discards the outcome $a_{i}$ and substitutes it with a random outcome for every party $i$ whose setting $x_{i}\neq\xi_{i}$. If a conditional probability distribution $(p(\vec{a}\mid\vec{x}))_{\vec{a},\vec{x}}$ belongs to the quantum ICO set, then $\Pi_{\vec{\xi}}(p)$ can be realized by a process matrix along with the canonical instruments. This can be formulated through the following SDP:
find $S$ (29)
subject to $\forall\vec{a}~{}\forall\vec{x}\quad\Pi_{\vec{\xi}}(p)(\vec{a}\mid\vec{x})=\operatorname{Tr}(SP_{\vec{\xi},\vec{a},\vec{x}})$
$S\in\operatorname{DualAff}(\operatorname{Choi}(\mathsf{NoSig}))$
$S\geq 0\,.$
We have in total $\prod_{i}n_{i}$ ($n_{i}$ is the number of settings for the $i$-th party) necessary conditions in the form of the SDP (29), corresponding to the choices of triggers.
Appendix B Biased OCB correlations
To show that the value $(1+\alpha+\sqrt{1+\alpha^{2}})/2$ is maximal for a biased OCB correlation, we decompose the biased OCB correlation into a random mixture of two single-trigger correlations, i.e.
$P(a_{1}=b\mid c=0)+\alpha\,P(a_{2}=x_{1}\mid c=1)$
$=\sum_{z\in\{0,1\}}\frac{1}{2}\Big{(}P(a_{1}=b\mid x_{1}=z,c=0)$
$\qquad+\alpha\,P(a_{2}=x_{1}\mid b=z,c=1)\Big{)}\,.$ (30)
The proof is then complete with the following lemma, which indicates that both correlations in the decomposition is upper bounded by $(1+\alpha+\sqrt{1+\alpha^{2}})/2$.
Lemma 1
For any $x_{*},b_{*}\in\{0,1\}$ and $\alpha\in\mathbb{R}$, the ICO bound of the following single-trigger correlation
$P(a_{1}=b\mid x_{1}=x_{*},c=0)$
$+\alpha\,P(a_{2}=x_{1}\mid b=b_{*},c=1)$ (31)
is $\left(1+\alpha+\sqrt{1+\alpha^{2}}\right)/2$.
Proof. To derive an upper bound of the single-trigger correlation (31), we construct a feasible solution of the dual SDP (23). We claim that the following positive operator $C$ on a 7-qubit Hilbert space $(\mathcal{H}_{1}\otimes\mathcal{H}_{2,3})\otimes(\mathcal{H}_{4}\otimes\mathcal{H}_{5,6,7})$ is equal to $\left((1+\alpha)+\sqrt{1+\alpha^{2}}\right)/2$ times the Choi operator of a no-signalling channel
$C:=\Omega+\sum_{k,l}|\Psi_{k,l}\rangle\langle\Psi_{k,l}|\,,$ (32)
where $\Omega$ is the single-trigger operator of the correlation (31):
$\begin{split}\Omega&:=\frac{1}{2}|00\rangle\langle 00|\otimes|x_{*}\rangle\langle x_{*}|\otimes|I\rangle\!\rangle\langle\!\langle I|\otimes|0\rangle\langle 0|\otimes|0\rangle\langle 0|\\ &\quad+\frac{1}{2}|11\rangle\langle 11|\otimes|x_{*}\rangle\langle x_{*}|\otimes|I\rangle\!\rangle\langle\!\langle I|\otimes|1\rangle\langle 1|\otimes|0\rangle\langle 0|\\ &\quad+\frac{\alpha}{2}|I\rangle\!\rangle\langle\!\langle I|\otimes|0\rangle\langle 0|\otimes|00\rangle\langle 00|\otimes|b_{*}\rangle\langle b_{*}|\otimes|1\rangle\langle 1|\\ &\quad+\frac{\alpha}{2}|I\rangle\!\rangle\langle\!\langle I|\otimes|1\rangle\langle 1|\otimes|11\rangle\langle 11|\otimes|b_{*}\rangle\langle b_{*}|\otimes|1\rangle\langle 1|\\ \end{split}$ (33)
and
$|\Psi_{k,l}\rangle=\sum_{i,j}c_{i,j,k,l}\,|i,i\rangle\otimes|k\rangle\otimes|j,j\rangle\otimes|l,0\rangle\,,$ (34)
with real coefficients $c_{i,j,k,l}$. The conditions that $C$ is proportional to the Choi operator of a no-signalling channel are
$\operatorname{Tr}_{2,3}C=\frac{I_{1}}{2}\otimes\operatorname{Tr}_{1,2,3}C\,,\quad\operatorname{Tr}_{5,6,7}C=\frac{I_{4}}{2}\otimes\operatorname{Tr}_{4,5,6,7}C\,,$ (35)
which are equivalent to the following equations:
$\forall j,l\quad$ $\sum_{k}c_{1,j,k,l}^{2}-c_{0,j,k,l}^{2}=\frac{(-1)^{l}}{2}\,,$ (36)
$\forall i,k\quad$ $\sum_{l}c_{i,1,k,l}^{2}-c_{i,0,k,l}^{2}=\frac{(-1)^{k}\alpha}{2}\,,$ (37)
$\forall l\quad$ $\sum_{k}c_{1,0,k,l}\,c_{1,1,k,l}-c_{0,0,k,l}\,c_{0,1,k,l}=\frac{(-1)^{l}}{2}\,,$ (38)
$\forall k\quad$ $\sum_{l}c_{0,1,k,l}\,c_{1,1,k,l}-c_{0,0,k,l}\,c_{1,0,k,l}=\frac{(-1)^{k}\alpha}{2}\,.$ (39)
We find the following solution of Eq. (36)-(39)
$c_{i,j,k,l}=\frac{\left(1+\alpha^{2}\right)^{1/4}}{4}\left(1-\frac{(-1)^{i+l}}{\sqrt{1+\alpha^{2}}}-\frac{(-1)^{j+k}\alpha}{\sqrt{1+\alpha^{2}}}\right)\,,$ (40)
with which $\operatorname{Tr}(C)=\operatorname{Tr}(\Omega)+\sum_{i,j,k,l}|c_{i,j,k,l}|^{2}=2(1+\alpha)+2\sqrt{1+\alpha^{2}}$, which is $\left((1+\alpha)+\sqrt{1+\alpha^{2}}\right)/2$ times the trace of the Choi operator of a no-signalling channel. Therefore, $\left((1+\alpha)+\sqrt{1+\alpha^{2}}\right)/2$ is an upper bound of the single-trigger correlation (31).
The bound can be attained with the instruments and a modification of the process matrix proposed in Ref. Oreshkov et al. (2012). Suppose that Alice and Bob share the following process matrix
$\begin{split}S_{{\rm OCB},\alpha}&=\frac{1}{4}I^{\otimes 4}+\frac{\alpha}{4\sqrt{1+\alpha^{2}}}I\otimes Z\otimes Z\otimes I\\ &\quad+\frac{1}{4\sqrt{1+\alpha^{2}}}Z\otimes I\otimes X\otimes Z\,,\end{split}$ (41)
where $X$ and $Z$ are Pauli-$X$ and Pauli-$Z$ matrices, respectively. To attain the bound, Alice measures her input qubit in computational basis and reprepare the her setting $x_{1}$ in the same basis, while Bob
  • (in the case $c=1$) measures his qubit in computational basis and reprepares an arbitrary qubit state;
  • (in the case $c=0$) measures his qubit in Fourier basis and reprepare $a_{2}\oplus b$ in computation basis.
The Choi representation of Alice’s and Bob’s instruments are
$M_{a_{1}\mid x_{1}}=|a_{1}\rangle\langle a_{1}|\otimes|x_{1}\rangle\langle x_{1}|\,,$ (42)
and
$N_{a_{2}\mid b,0}$ $=|X_{a_{2}}\rangle\langle X_{a_{2}}|\otimes|a_{2}\oplus b\rangle\langle a_{2}\oplus b|\,,$ (43)
$N_{a_{2}\mid b,1}$ $=|a_{2}\rangle\langle a_{2}|\otimes\rho\,,$ (44)
where $\rho$ is an arbitrary density operator on a qubit and $|X_{a_{2}}\rangle=(|0\rangle+(-1)^{a_{2}}|1\rangle)/\sqrt{2}$ are the eigenstates of Pauli-$X$ matrix (Fourier basis). The joint conditional probability distribution between Alice’s and Bob’s outcomes $p(a_{1},a_{2}\mid x_{1},b,c)=\operatorname{Tr}\left[{S_{\rm OCB,\alpha}}^{T}(M_{a_{1}\mid x_{1}}\otimes N_{a_{2}\mid b,c})\right]$ is computed to be
$p(a_{1},a_{2}\mid x_{1},b,0)$ $=\frac{1}{4}+\frac{(-1)^{a_{1}+b}}{4\sqrt{1+\alpha^{2}}}\,,$ (45)
$p(a_{1},a_{2}\mid x_{1},b,1)$ $=\frac{1}{4}+\frac{(-1)^{a_{2}+x_{1}}\,\alpha}{4\sqrt{1+\alpha^{2}}}\,.$ (46)
It follows that the correlation achieves the bound:
$P(a_{1}=b\mid x_{1}=x_{*},c=0)+\alpha\,P(a_{2}=x_{1}\mid b=b_{*},c=1)$
$=\frac{1}{2}\sum_{a_{1},a_{2},x_{1},b}\delta_{a_{1},b}\,\delta_{x_{1},x_{*}}\,p(a_{1},a_{2}\mid x_{1},b,0)$
$\quad+\frac{\alpha}{2}\sum_{a_{1},a_{2},x_{1},b}\delta_{a_{2},x_{1}}\,\delta_{b,b_{*}}\,p(a_{1},a_{2}\mid x_{1},b,1)$
$=\frac{1}{2}\left(1+\frac{1}{\sqrt{1+\alpha^{2}}}\right)+\frac{\alpha}{2}\left(1+\frac{\alpha}{\sqrt{1+\alpha^{2}}}\right)$
$=\frac{1}{2}\left(1+\alpha+\sqrt{1+\alpha^{2}}\right)\,.$ (47)
$\blacksquare$
Appendix C Biased LGYNI correlations
Figure 2: Visualizing correlations and probability distributions. (a) Causal bound, ICO bound and algebraic maximum of the success probability of biased LGYNI. (b) Geometry of probability distributions on a 2-dimensional plane. The green area represents the projection of causal probability distributions; the blue area represents the projection of probability distributions exhibiting ICO. The remaining area (violet area) represents the non-causal probability distributions beyond the ICO ones.
We now introduce a biased version of the LGYNI game, corresponding to the correlation
$\mathcal{I}_{{\rm LGYNI},\alpha}=\alpha\,P_{\rm succ}^{11}+(1-\alpha)\,P_{\rm succ}^{\rm ave}\,,$ (48)
where $\alpha$ ranges between 0 and 1 and $P_{\rm succ}^{\rm ave}:=\frac{P_{\rm succ}^{01}+P_{\rm succ}^{10}}{2}$ is the average of $P_{\rm succ}^{01}$ and $P_{\rm succ}^{10}$. Clearly, the correlation $\mathcal{I}_{{\rm LGYNI},\alpha}$ is single-trigger, and for $\alpha=1/3$ its maximization is equivalent to the maximization of the probability of success in the original LGYNI game.
For the biased LGYNI game, one has the causal inequality
$\mathcal{I}_{{\rm LGYNI},\alpha}\leq\max\left\{1-\alpha,\frac{1+\alpha}{2}\right\}\,.$ (49)
The proof is as follows: let $p_{A}$ and $p_{B}$ be the marginal probability distributions of Alice’s and Bob’s outcomes, defined as
$p_{A}(a_{1}\mid x_{1},x_{2})$ $:=\sum_{a_{2}}p(a_{1},a_{2}\mid x_{1},x_{2})\,,$ (50)
$p_{B}(a_{2}\mid x_{1},x_{2})$ $:=\sum_{a_{1}}p(a_{1},a_{2}\mid x_{1},x_{2})$ (51)
respectively. If Alice’s operation causally precedes Bob’s, then
$p(1,1\mid 1,1)\leq p_{A}(1\mid 1,1)=p_{A}(1\mid 1,0)$ (52)
In this case, biased LGYNI is upper bounded as follows
$\mathcal{I}_{{\rm LGYNI},\alpha}$
$=\alpha\,P_{\rm succ}^{11}+(1-\alpha)\,\frac{P_{\rm succ}^{01}+P_{\rm succ}^{10}}{2}$
$=\alpha p(1,1\mid 1,1)+\frac{1-\alpha}{2}p_{A}(0\mid 1,0)+\frac{1-\alpha}{2}p_{B}(0\mid 0,1)$
$\leq\alpha p_{A}(1\mid 1,0)+\frac{1-\alpha}{2}p_{A}(0\mid 1,0)+\frac{1-\alpha}{2}p_{B}(0\mid 0,1)$
$=\alpha(1-p_{A}(0\mid 1,0))+\frac{1-\alpha}{2}p_{A}(0\mid 1,0)$
$\quad+\frac{1-\alpha}{2}p_{B}(0\mid 0,1)$
$\leq\max\left\{\alpha,\frac{1-\alpha}{2}\right\}+\frac{1-\alpha}{2}$
$=\max\left\{\frac{1+\alpha}{2},1-\alpha\right\}\,,$ (53)
where the first inequality is due to (52), and the second inequality is due to the monotonicity as a linear function of $p_{A}(0\mid 1,0)$ and the probability constraint $p_{B}\leq 1$. Vice versa, biased LGYNI is upper bounded by the same value when Bob’s operation precedes Alice’s. Since a random mixture can not increase the bound, the causal bound of biased LGYNI is no greater than (53). The value (53) can be saturated by the following causal process: Alice sends her classical bit $x$ to Bob; she produces outcome 0 if $\alpha<\frac{1-\alpha}{2}$ and produces outcome 1 otherwise; Bob produces his outcome to be $x$.
Quantum processes with ICO can violate the inequality (49) for suitable values of the parameter $\alpha$. Explicitly, the ICO bound for the biased LGYNI game can be computed as an SDP. The result is shown in Figure 2(a) for all possible values of $\alpha$ in the interval $[0,1]$. Interestingly, it appears that the biased LGYNI causal inequality cannot be violated by any quantum ICO process for $\alpha\leq 0.188$. (In the last part of this section, we provide a rigorous proof of this fact for $\alpha\leq\frac{4-\sqrt{5}}{11}\approx 0.16$. ) To the best of our knowledge, biased LGYNI with $0<\alpha\leq 0.188$ is the first non-trivial example of a causal inequality with no quantum violation in the ICO framework.
At first sight, the inequality (49) for $\alpha\leq 0.188$ may appear as a causal analogue of Bell inequalities with no quantum violations Almeida et al. (2010); Fritz et al. (2013). A more in-depth analysis of the geometry of ICO correlations, however, reveals that this inequality is not tight, meaning that it does not identify a facet of the quantum ICO set.
Let us visualize the success probabilities $P_{\rm succ}^{11}$ and $P_{\rm succ}^{\rm ave}$ in a two-dimensional plane. The values of $P_{\rm succ}^{11}$ and $P_{\rm succ}^{\rm ave}$ compatible with the causal inequality (49) and with the ICO bound are shown in Figure 2(b), where they correspond to the green area and the blue area, respectively. The constraints arising from the causal inequality can be equivalently characterized by the condition
$P_{\rm succ}^{11}+2P_{\rm succ}^{\rm ave}\leq 2\,,\$ (54)
The condition (54) is the causal inequality associated with the canonical LGYNI game, which corresponds to a facet of the causal polytope Branciard et al. (2015).
The shape of the boundary of the quantum ICO set, instead, shows that the set of probability distributions generated by quantum ICO processes is not a polytope, i.e. it has infinite extreme points. Figure 2(b) also shows that the boundary of the quantum ICO set is distinct from the boundary of the causal polytope for every value except for the extreme points of the causal polytope, corresponding to the values $p_{\rm succ}^{11}=0$ and $p_{\rm succ}^{11}=1$. The fact that the quantum ICO and causal boundaries do not coincide implies that the causal inequality (49) is not tight. An interesting open question is whether there exist examples of tight causal inequalities with no quantum ICO violation, namely whether there exists flat faces of the quantum set that coincide with faces of the causal polytope.
Now we rigorously prove that when $\alpha\in\left[0,\frac{4-\sqrt{5}}{11}\right]$, the ICO bound of biased LGYNI is equal to its causal bound $1-\alpha$. We claim that the follow positive operator $C$ on a 6-qubit Hilbert space $(\mathcal{H}_{1}\otimes\mathcal{H}_{2,3})\otimes(\mathcal{H}_{4}\otimes\mathcal{H}_{5,6})$ is equal to $(1-\alpha)$ times the Choi operator of a no-signalling channel
$C:=\Omega+|\Psi_{01}\rangle\langle\Psi_{01}|+|\Psi_{10}\rangle\langle\Psi_{10}|+|\Psi_{00}\rangle\langle\Psi_{00}|\,,$ (55)
where $\Omega$ is the single-trigger operator of biased LGYNI
$\Omega$ $:=\,\alpha|11\rangle\langle 11|\otimes|1\rangle\langle 1|\otimes|11\rangle\langle 11|\otimes|1\rangle\langle 1|$
$\,+\frac{1-\alpha}{2}\,|00\rangle\langle 00|\otimes|1\rangle\langle 1|\otimes|I\rangle\!\rangle\langle\!\langle I|\otimes|0\rangle\langle 0|$
$\,+\frac{1-\alpha}{2}\,|I\rangle\!\rangle\langle\!\langle I|\otimes|0\rangle\langle 0|\otimes|00\rangle\langle 00|\otimes|1\rangle\langle 1|\,,$ (56)
and
$|\Psi_{01}\rangle$ $:=c_{0}|11\rangle\otimes|0\rangle\otimes|11\rangle\otimes|1\rangle$ (57)
$\quad+c_{1}|00\rangle\otimes|0\rangle\otimes|11\rangle\otimes|1\rangle$
$|\Psi_{10}\rangle$ $:=c_{0}|11\rangle\otimes|1\rangle\otimes|11\rangle\otimes|0\rangle$ (58)
$\quad+c_{1}|11\rangle\otimes|1\rangle\otimes|00\rangle\otimes|0\rangle$
$|\Psi_{00}\rangle$ $:=c_{2}|11\rangle\otimes|0\rangle\otimes|11\rangle\otimes|0\rangle$ (59)
$\quad+c_{3}|00\rangle\otimes|0\rangle\otimes|11\rangle\otimes|0\rangle$
$\quad+c_{3}|11\rangle\otimes|0\rangle\otimes|00\rangle\otimes|0\rangle\,,$
with non-negative coefficients $c_{0},\cdots,c_{3}$. The conditions that $C$ is proportional to the Choi operator of a no-signalling channel are
$\operatorname{Tr}_{2,3}C=\frac{I_{1}}{2}\otimes\operatorname{Tr}_{1,2,3}C\,,\quad\operatorname{Tr}_{5,6}C=\frac{I_{4}}{2}\otimes\operatorname{Tr}_{4,5,6}C\,,$ (60)
which is equivalent to the following equations:
$c_{1}^{2}-c_{0}^{2}$ $=\alpha\,,$ (61)
$c_{0}^{2}+c_{2}^{2}-c_{3}^{2}$ $=\frac{1-\alpha}{2}\,,$ (62)
$c_{1}^{2}+c_{3}^{2}$ $=\frac{1-\alpha}{2}\,.$ (63)
$c_{0}c_{1}+c_{2}c_{3}$ $=\frac{1-\alpha}{2}\,,$ (64)
We can solve $c_{1}$, $c_{2}$ and $c_{3}$ in terms of $c_{0}$ according to Eqs. (61)-(63)
$c_{1}$ $=\sqrt{c_{0}^{2}+\alpha}\,,$ (65)
$c_{2}$ $=\sqrt{1-2\alpha-2c_{0}^{2}}\,,$ (66)
$c_{3}$ $=\sqrt{\frac{1-3\alpha}{2}-c_{0}^{2}}\,.$ (67)
It follows that
$\frac{\operatorname{Tr}(C)}{4}=\frac{\alpha+2(1-\alpha)+2\left(c_{0}^{2}+c_{1}^{2}\right)+\left(c_{2}^{2}+2c_{3}^{2}\right)}{4}=1-\alpha$ (68)
according to Eqs. (65)-(67). Now it suffices to show the existence of $c_{0},\cdots,c_{3}$ satisfying Eqs. (61)-(64). With the solutions in Eqs. (65)-(67), we focus on the quantity $f(c_{0})=c_{0}c_{1}+c_{2}c_{3}-\frac{1-\alpha}{2}$, corresponding to Eq. (64). When $\alpha\in\left[0,\frac{4-\sqrt{5}}{11}\right]$, it holds that
$f(0)=\sqrt{\frac{(1-2\alpha)(1-3\alpha)}{2}}-\frac{1-\alpha}{2}\geq 0\,;$ (69)
and
$f\left(\sqrt{\frac{1-3\alpha}{2}}\right)=\sqrt{\frac{1-3\alpha}{2}}\sqrt{\frac{1-\alpha}{2}}-\frac{1-\alpha}{2}\leq 0\,;$ (70)
By continuity of the function $f$, there exists $c_{0}\in\left[0,\sqrt{\frac{1-3\alpha}{2}}\right]$ such that $f(c_{0})=0$ and thus Eq. (64) is satisfied. Therefore, $(1-\alpha)$ is an upper bound of $\eta(\Omega)$ and thus an upper bound of the biased LGYNI. This completes the proof.
Appendix D Characterization of supermaps in bistochastic quantum and classical theories
Figure 3: Circuit realization of simplest classical supermaps with definite causal order. (a) no-signalling supermaps ($A\not\preceq B$ and $B\not\preceq A$), (b) unidirectional signalling supermaps ($A\preceq B$ or $B\preceq A$) with fixed time direction, (c) unidirectional signalling supermaps ($A\preceq B$ or $B\preceq A$) with dynamical time direction. In this figure, the blue half discs correspond to classical bits, and the blue boxes correspond to transformations that are either identity or bit-flip.
In this section, we show the proof of the characterization of quantum supermaps transforming a pair of bistochastic channels. We then provide a classification of the set of simplest bipartite classical supermaps, considering the classical limits of bistochastic quantum theory. This set includes the process used to achieve a guaranteed win in the GYNI game, as discussed in the main text.
Characterization of quantum supermaps
We consider a supermap transforming a pair of bistochastic channels, one mapping states on system $A_{\rm I}$ to to states on system $A_{\rm O}$ and the other mapping states on system $B_{\rm I}$ to states on system $B_{\rm O}$. The output of the supermap is a channel from system $C_{\rm I}$ to system $C_{\rm O}$. Since a supermap is a completely positive map, its Choi operator $S$ has to be a positive operator. The requirement that $S$ transforms any pair of bistochastic channels into an ordinary channel is equivalent to the normalization condition
$\operatorname{Tr}[(\rho_{C_{\rm I}}\otimes I_{C_{\rm O}}\otimes M_{A_{\rm I}A_{\rm O}}\otimes N_{B_{\rm I}B_{\rm O}})^{T}S]=1\,,$ (71)
where $\rho$ is an arbitrary density operator on system $C_{\rm I}$, and $M$ ($N$) is the Choi operator of an arbitrary bistochastic channel from $A_{\rm I}$ ($B_{\rm I}$) to $A_{\rm O}$ ($B_{\rm O}$). Specifically, as the Choi operator $M$ can be written as Chiribella and Liu (2022)
$M=\frac{I_{A_{\rm I}}\otimes I_{A_{\rm O}}}{d_{A}}+T$ (72)
where $T$ is an operator such that $\operatorname{Tr}_{A_{\rm I}}T=0$ and $\operatorname{Tr}_{A_{\rm O}}T=0$, the subspace spanned by $M$ is characterized by the projection $\Pi_{A}$
$\Pi_{A}(M)$ $:={}_{[A_{\rm I}A_{\rm O}]}M+{}_{[1-A_{\rm I}-A_{\rm O}+A_{\rm I}A_{\rm O}]}M$
$={}_{[1-A_{\rm I}-A_{\rm O}+2A_{\rm I}A_{\rm O}]}M\,,$ (73)
where we have used the notation ${}_{[X]}S:=\operatorname{Tr}_{X}[S]\otimes\frac{I_{X}}{d_{X}}$ introduced in Methods. Similarly, the subspace spanned by $N$ is characterized by the projection $\Pi_{B}$
$\Pi_{B}(N):={}_{[1-B_{\rm I}-B_{\rm O}+2B_{\rm I}B_{\rm O}]}N\,.$ (74)
Let $S^{\prime}$ be the the reduced operator $S^{\prime}:=\operatorname{Tr}_{C_{\rm O}}S$. It follows that condition (71) is equivalent to
$\Pi_{A}\otimes\Pi_{B}(S^{\prime})=\frac{I_{A_{\rm I}A_{\rm O}B_{\rm I}B_{\rm O}C_{\rm I}}}{d_{A}d_{B}}\,,$ (75)
which is in turn equivalent to the following two conditions
$\operatorname{Tr}(S)=d_{A}d_{B}d_{C_{\rm I}}\,,$ (76)
$\Pi_{A}\otimes\Pi_{B}(S^{\prime})={}_{[A_{\rm I}A_{\rm O}B_{\rm I}B_{\rm O}C_{\rm I}]}S^{\prime}\,.$ (77)
Eq. (77) can be decomposed into four constraints
${}_{[A_{\rm I}A_{\rm O}B_{\rm I}B_{\rm O}(1-C_{\rm I})C_{\rm O}]}S=0\,,$ (78)
${}_{[(1-A_{\rm I})(1-A_{\rm O})B_{\rm I}B_{\rm O}C_{\rm O}]}S=0\,,$
${}_{[(1-B_{\rm I})(1-B_{\rm O})A_{\rm I}A_{\rm O}C_{\rm O}]}S=0\,,$
${}_{[(1-A_{\rm I})(1-A_{\rm O})(1-B_{\rm I})(1-B_{\rm O})C_{\rm O}]}S=0\,,$
which can be regarded as commuting projections. This completes the proof. The characterization presented in Methods is a special case of (76) and (78) with trivial systems $C_{\rm I}$ and $C_{\rm O}$.
Simplest bipartite classical supermaps
In the simplest case, the systems $A_{\rm I}$, $A_{\rm O}$, $B_{\rm I}$ and $B_{\rm O}$ are all qubits, while the systems $C_{\rm I}$ and $C_{\rm O}$ are trivial. Classical supermaps can be represented by operators diagonal in the computational basis of $\mathcal{H}_{A_{\rm I}}\otimes\mathcal{H}_{A_{\rm O}}\otimes\mathcal{H}_{B_{\rm I}}\otimes\mathcal{H}_{B_{\rm O}}$
$S=\sum_{i,j,k,l\in\{0,1\}}S_{i,j,k,l}\,|i\rangle\langle i|\otimes|j\rangle\langle j|\otimes|k\rangle\langle k|\otimes|l\rangle\langle l|\,.$ (79)
Theorem 4 in Methods implies that the coefficients in Eq. (79) are probabilities and satisfy
$\begin{split}\sum_{i=j,\,k=l}S_{i,j,k,l}&=1\,,\\ \sum_{i=j,\,k=l\oplus 1}S_{i,j,k,l}&=1\,,\\ \sum_{i=j\oplus 1,\,k=l}S_{i,j,k,l}&=1\,,\\ \sum_{i=j\oplus 1,\,k=l\oplus 1}S_{i,j,k,l}&=1\,.\end{split}$ (80)
It follows that these classical supermaps form a convex set with $2^{8}$ extreme points:
$\sum_{m=0}^{3}|i_{m}\rangle\langle i_{m}|\otimes|j_{m}\rangle\langle j_{m}|\otimes|k_{m}\rangle\langle k_{m}|\otimes|l_{m}\rangle\langle l_{m}|$ (81)
where
$i_{0}=j_{0},\,k_{0}=l_{0}\,,$ (82)
$i_{1}=j_{1},\,k_{1}=l_{1}\oplus 1\,,$ (83)
$i_{2}=j_{2}\oplus 1,\,k_{2}=l_{2}\,,$ (84)
$i_{3}=j_{3}\oplus 1,\,k_{3}=l_{3}\oplus 1\,.$ (85)
We provide a classification of the $2^{8}$ extreme points, corresponding to deterministic classical supermaps without a pre-defined direciton of time:
  • No-signalling supermaps ($A\not\preceq B$ and $B\not\preceq A$). There are $2^{4}$ points of this type, all equivalent to the circuit in Figure 3(a), up to interchange of input/output systems;
  • Unidirectional signalling supermaps ($A\preceq B$ or $B\preceq A$) with fixed time direction. There are $2^{5}$ points of this type, all equivalent to the circuit in Figure 3(b), up to interchange of input/output systems and causal orders;
  • Unidirectional signalling supermaps ($A\preceq B$ or $B\preceq A$) with dynamically controlled time direction. There are $2^{6}$ points of this type, all equivalent to the circuit in Figure 3(c) where the time direction of $B$ is dynamically decided by the output of $A$, up to interchange of input/output systems and causal orders;
  • Supermaps with indefinite causal order and time direction. There are $2^{4}+2^{7}$ points of this type. Up to interchange of input/output systems, $2^{4}$ points of them can be expressed through functions: $i_{m}=j_{m}\oplus l_{m}\oplus\alpha$ and $k_{m}=j_{m}\oplus l_{m}\oplus\beta$ where $\alpha,\beta\in\{0,1\}$. The other $2^{7}$ points, however, cannot be understood as stochastic processes.
The classical supermap used to achieve a guaranteed win in the GYNI game in the main text corresponds to the extreme point given by:
$S_{0,0,0,0}=S_{1,1,1,0}=S_{1,0,1,1}=S_{0,1,0,1}=1\,.$ (86)
Appendix E Reaching the algebraic maximum in certain scenarios
In this section, we provide constructions of strategies in bistochastic classical theory which realize algebraically maximal correlations. Let $(N,n,m)$ denote the scenario with $N$ parties, each with $n$ settings and $m$ outcomes. We show that perfect two-way signalling of ternary digits can be achieved with choices of classical supermaps and local bistochastic instruments. Equipped with a perfect two-way signalling of each other’s setting, the two parties can generate arbitrary deterministic conditional probability distribution and thus the BICO value in the (2,3,?) scenario is equal to the algebraic maximum. Moreover, we show that if there are only 2 settings for each party (the (2,2,?) scenario), every bipartite probability distribution can be generated with choices of classical supermaps and local bistochastic instruments. Our constructions cannot be directly extended to the scenarios with more than 3 settings. This leads us to raise the open question whether there are non-trivial constraints on the correlations generated in bistochastic classical and quantum theory in general scenarios.
Let us consider bipartite classical supermaps transforming a pair of classical bistochastic instruments to probabilities. These supermaps can be represented by operators $S$ diagonal in an orthonormal basis on the Hilbert space $\mathcal{H}_{A_{\rm I}}\otimes\mathcal{H}_{A_{\rm O}}\otimes\mathcal{H}_{B_{\rm I}}\otimes\mathcal{H}_{B_{\rm O}}$ corresponding to Alice’s and Bob’s input/output systems
$S=\sum_{i,j,k,l\in\{0,...,d-1\}}S_{i,j,k,l}\,|i\rangle\langle i|\otimes|j\rangle\langle j|\otimes|k\rangle\langle k|\otimes|l\rangle\langle l|\,$ (87)
where $d$ is the dimension of every input/output system and $S_{i,j,k,l}$ are probabilities.
Perfect two-way signalling in the (2,3,?) scenerio
In the scenario where each of the two parties has $n\leq 3$ settings, there is a strategy generalizing the one introduced in the main text for winning the GYNI game, which realizes two-way signalling of $n$-nary digits. In this strategy, Alice uses the classical instrument
$q^{(1)}_{a_{1}\mid x_{1}}(s^{\prime}_{1}\mid s_{1})=\delta_{a_{1},s^{\prime}_{1}}\delta_{s_{1}^{\prime},((s_{1}-x_{1})\mod n)}\,,$
and Bob uses the classical instrument
$q^{(2)}_{a_{2}\mid x_{2}}(s^{\prime}_{2}\mid s_{2})=\delta_{a_{2},s^{\prime}_{2}}\delta_{s_{2}^{\prime},((x_{2}-s_{2})\mod n)}\,.$
Alice’s and Bob’s local experiments are connected by a deterministic classical supermap, represented as a function that maps the final values of the classical systems $(s^{\prime}_{1},s^{\prime}_{2})$ into their initial values $(s_{1},s_{2})$ according to the rule
$s_{1}=((s^{\prime}_{1}+s^{\prime}_{2})\mod n)\,,\quad s_{2}=((s^{\prime}_{1}-s^{\prime}_{2})\mod n)\,.$ (88)
[…]
The (2,2,?) scenario
If each party has 2 settings, we can construct a class of classical supermaps which not only enable perfect signalling of bits but also distribute shared randomness between the two parties. This allows the two parties to create any random mixture of deterministic conditional probability distributions. The class of supermaps, denoted by $\{\mathcal{S}_{x,y}\}$, can be represented by positive operators $\{S_{x,y}\}$ on the Hilbert space $\mathcal{H}_{A_{\rm O}}\otimes\mathcal{H}_{B_{\rm O}}\otimes\mathcal{H}_{A_{\rm I}}\otimes\mathcal{H}_{B_{\rm I}}$:
$S_{x,y}$ $:=(|x,x\rangle\langle x,x|+|y,y\rangle\langle y,y|)\otimes(I-|y\rangle\langle y|)\otimes|x\rangle\langle x|$
$+(|x,y\rangle\langle x,y|+|y,x\rangle\langle y,x|)\otimes|y\rangle\langle y|\otimes(I-|x\rangle\langle x|)$
$+(I-|x\rangle\langle x|-|y\rangle\langle y|)^{\otimes 2}\otimes|y\rangle\langle y|\otimes|x\rangle\langle x|$ (89)
where $x,y\in\{0,\cdots,d-1\}$. We can check that the positive operators (89) satisfy the conditions (76) and (78). Assuming that $d$ is even, we construct the classical bistochastic instrument of party $i$ to be a deterministic function $q^{(i)}_{a_{i}\mid x_{i}}(s^{\prime}_{i}\mid s_{i})$ which gives the outcome $a_{i}=f_{i}^{(t)}(x_{i},z)$ and produces the classical state $s^{\prime}_{i}=2t+z$, with $t=\lfloor\frac{s_{i}}{2}\rfloor$, $z=((x_{i}+s_{i})\mod 2)$ and $f_{i}^{(t)}$ is an function depending on the parameter $t\in\{0,\cdots,\frac{d}{2}-1\}$. If the local experiments are connected by the supermap $\mathcal{S}_{2t,2t+1}$, then the correlation between the two parties is the following deterministic probability distribution
$p(a_{1},a_{2}\mid x_{1},x_{2})=\delta_{a_{1},f_{1}^{(t)}(x_{1},x_{2})}\delta_{a_{2},f_{2}^{(t)}(x_{2},x_{1})}\,.$ (90)
An arbitrary random mixture of deterministic probability distributions can be generated by randomly choosing the supermap from $\{\mathcal{S}_{2t,2t+1}\}$. In other words, the above strategy can be used to generate arbitrary bipartite probability distribution in the (2,2,?) scenario.
Appendix F Proof of Theorem 1 in Methods
Consider the realization of the instruments $\mathcal{M}_{a_{i}\mid x_{i}}^{(i)}$ by an isometry and a projective measurement. Let $\left\{K_{a_{i}\mid x_{i},j}^{(i)}\right\}$ be a Kraus decomposition of the instrument $\mathcal{M}_{a_{i}\mid x_{i}}^{(i)}$ of the $i$-th party. For every classical label $x_{i}$, define $V_{x_{i}}^{(i)}$ to be the following isometry from $\mathcal{H}_{A_{\rm I}^{(i)}}$ to $\mathcal{H}_{L}\otimes\mathcal{H}_{A_{\rm O}^{(i)}}\otimes\mathcal{H}_{J}$
$V^{(i)}_{x_{i}}:=\sum_{a_{i},j}|a_{i}\rangle_{L}\otimes K_{a_{i}\mid x_{i},j}^{(i)}\otimes|j\rangle_{J}\,,$ (91)
where $L$ is the classical register for the outcome $a_{i}$ and $J$ the classical register for the index $j$. Choosing the associated Hilbert spaces to be $\mathcal{H}_{L}\simeq\mathbb{C}^{m_{i}}$ and $\mathcal{H}_{J}\simeq\mathcal{H}_{A_{\rm I}^{(i)}}\otimes\mathcal{H}_{A_{\rm O}^{(i)}}$, the two systems $L$ and $J$ have no dependence on the setting $x_{i}$. Extending $V_{x_{i}}^{(i)}$ to an unitary $U_{x_{i}}^{(i)}$ from $\mathcal{H}_{A_{\rm I}^{(i)}}\otimes\mathcal{H}_{E_{\rm I}^{(i)}}$ to $\mathcal{H}_{L}\otimes\mathcal{H}_{A_{\rm O}^{(i)}}\otimes\mathcal{H}_{J}\otimes\mathcal{H}_{E_{\rm O}^{(i)}}$ with some auxiliary systems $E_{\rm I}^{(i)}$ and $E_{\rm O}^{(i)}$, the map $\mathcal{M}_{a_{i}\mid x_{i}}^{(i)}$ can be realized as
$\rho\mapsto\operatorname{Tr}_{L,J,E_{\rm O}^{(i)}}\left[\Pi_{a_{i}}U_{x_{i}}^{(i)}\left(\rho\otimes|0\rangle\langle 0|_{E_{\rm I}^{(i)}}\right){U_{x_{i}}^{(i)}}^{\dagger}\Pi_{a_{i}}\right]\,,$ (92)
where
$\Pi_{a_{i}}=|a_{i}\rangle\langle a_{i}|_{L}\otimes I_{A_{\rm O}^{(i)}}\otimes I_{J}\otimes I_{E_{\rm O}^{(i)}}\,.$ (93)
Define local operations $\mathcal{E}^{(i)}$ and $\mathcal{D}^{(i)}$ to be the channels
$\forall\rho\quad\mathcal{E}^{(i)}(\rho)$ $=\rho\otimes|0\rangle\langle 0|_{E_{\rm I}^{(i)}}\,,$ (94)
$\forall\sigma\quad\mathcal{D}^{(i)}(\sigma)$ $=\operatorname{Tr}_{L,J,E_{\rm O}^{(i)},F}\mathcal{U}^{(i)}(\sigma)\,,$ (95)
where $\mathcal{U}^{(i)}$ is the controlled unitary channel
$\sum_{x_{i}}U_{x_{i}}^{(i)}\otimes|x_{i}\rangle\langle x_{i}|_{F}\,.$ (96)
with the control system denoted by $F$.
We construct the projectors of the labelled projective instruments $\mathcal{M}_{a_{i}\mid x_{i}}^{(i)^{\prime}}$ to be $\left\{{U_{x_{i}}^{(i)}}^{\dagger}\Pi_{a_{i}}U_{x_{i}}^{(i)}\right\}_{a_{i}}$ (whose rank is independent of $a_{i}$ and $x_{i}$), the label of $\mathcal{M}_{a_{i}\mid x_{i}}^{(i)^{\prime}}$ to be $x_{i}$, and the process $\mathcal{S}^{\prime}$ to be the composite of the local operations $\mathcal{E}^{(i)}$, $\mathcal{D}^{(i)}$ and the process $\mathcal{S}$, i.e.
$\mathcal{S}^{\prime}\left(\mathcal{M}^{(1)^{\prime}}_{a_{1}\mid x_{1}},\cdots,\mathcal{M}^{(N)^{\prime}}_{a_{N}\mid x_{N}}\right)$ (97)
$=\mathcal{S}\left(\mathcal{D}^{(1)}\circ\mathcal{M}^{(1)^{\prime}}_{a_{1}\mid x_{1}}\circ\mathcal{E}^{(1)},\cdots,\mathcal{D}^{(N)}\circ\mathcal{M}^{(N)^{\prime}}_{a_{N}\mid x_{N}}\circ\mathcal{E}^{(N)}\right)$
According to the realization in Eq. (92), the local operations $\mathcal{E}^{(i)}$ and $\mathcal{D}^{(i)}$ transform the labelled projective instruments $\mathcal{M}_{a_{i}\mid x_{i}}^{(i)^{\prime}}$ to $\mathcal{M}_{a_{i}\mid x_{i}}^{(i)}$. Hence $\mathcal{S}^{\prime}\left(\mathcal{M}^{(1)^{\prime}}_{a_{1}\mid x_{1}},\cdots,\mathcal{M}^{(N)^{\prime}}_{a_{N}\mid x_{N}}\right)$ reproduces the joint conditional probability $\mathcal{S}\left(\mathcal{M}^{(1)}_{a_{1}\mid x_{1}},\cdots,\mathcal{M}^{(N)}_{a_{N}\mid x_{N}}\right)$.
Appendix G Proof of Theorem 2 in Methods
The key of the proof is the fact that a projective measurement followed by the measurement outcome being discarded can be written as a uniform mixture of unitary channels. Specifically, let $\{P_{k}\}_{k=1}^{n}$ be the operators of a projective measurement. Proposition 4.6 of Watrous’ textbook Watrous (2018) implies that there exist $2^{n}$ unitary gates $\{U_{l}\}_{l=1}^{2^{n}}$ such that
$\forall\rho\quad\sum_{k=1}^{n}P_{k}\rho P_{k}=\frac{1}{2^{n}}\sum_{l=1}^{2^{n}}U_{l}\rho U_{l}^{\dagger}\,.$ (98)
Consider the map $\mathcal{M}_{\mathcal{I}}$ associated to a given single trigger correlation $\sum_{\vec{a},\vec{x}}\alpha_{\vec{a},\vec{x}}\,p(\vec{a}\mid\vec{x})$, with triggers $\vec{\xi}$ and the labelled projective instruments $\mathcal{M}_{a_{i}\mid x_{i}}^{(i)}(\cdot)=P_{a_{i},x_{i}}^{(i)}\cdot P_{a_{i},x_{i}}^{(i)}\otimes|x_{i}\rangle\langle x_{i}|$. We observe that the value of $\mathcal{M}_{\mathcal{I}}$ is unchanged when substituting $\mathcal{M}_{a_{i}\mid x_{i}}^{(i)}$ with $\frac{1}{m_{i}}\sum_{k=0}^{m_{i}-1}M_{k\mid x_{i}}^{(i)}$ for any party $i$ and any setting $x_{i}\neq\xi_{i}$. For example, let $x_{1}$ be a non-trigger setting of the first party, i.e. $x_{1}\neq\xi_{1}$. Indeed, since $\alpha_{\vec{a},\vec{x}}$ does not depends on $a_{1}$, we have
$\sum_{a_{1}=0}^{m_{1}-1}\alpha_{\vec{a},\vec{x}}\mathcal{M}_{a_{1}\mid x_{1}}^{(1)}$ $=\frac{1}{m_{1}}\sum_{a_{1}=0}^{m_{1}-1}\sum_{k=0}^{m_{1}-1}\alpha_{\vec{a},\vec{x}}\mathcal{M}_{k\mid x_{1}}^{(1)}$
$=\sum_{a_{1}=0}^{m_{1}-1}\alpha_{\vec{a},\vec{x}}\left(\frac{\sum_{k=0}^{m_{1}-1}\mathcal{M}_{k\mid x_{1}}^{(1)}}{m_{1}}\right)\,.$ (99)
According to the property in Eq. (98), for every party $i$, there exists a collection of unitary channels $\left\{\mathcal{U}_{x_{i}}^{(i,l)}\right\}_{l=1}^{2^{m_{i}}}$ such that
$\sum_{k=0}^{m_{i}-1}\mathcal{M}_{k\mid x_{i}}^{(i)}=\frac{1}{2^{m_{i}}}\sum_{l=1}^{2^{m_{i}}}\mathcal{U}_{x_{i}}^{(i,l)}\otimes|x_{i}\rangle\langle x_{i}|\,.$ (100)
Iterating the above substitutions for every party $i$ and every setting $x_{i}\neq\xi_{i}$, $\mathcal{M}_{\mathcal{I}}$ is readily in the form of a uniform mixture of the maps $(\mathcal{M}_{\mathcal{I},j})_{j}$ presented in Theorem 2. This completes the proof.
Appendix H Proof of ICO bound of single-trigger correlations
Theorem 1 and 2 in Methods imply that the optimization of a single-trigger correlation can be restricted without loss of generality to an optimization over the following instruments
$\mathcal{M}_{\,a_{i}|x_{i}}^{(i)}(\rho)=\begin{cases}P_{a_{i},\xi_{i}}^{(i)}\,\rho\,P_{a_{i},\xi_{i}}^{(i)}\otimes|\xi_{i}\rangle\langle\xi_{i}|&x_{i}=\xi_{i}\,,\\ \frac{1}{m_{i}}\,U_{x_{i}}^{(i)}\,\rho\,U_{x_{i}}^{(i){\dagger}}\otimes|x_{i}\rangle\langle x_{i}|&x_{i}\neq\xi_{i}\,.\end{cases}$ (101)
where for each party $i$, $\left(P_{a_{i},\xi_{i}}^{(i)}\right)_{a_{i}=0}^{m_{i}-1}$ are projectors of equal rank and $\left(U_{x_{i}}^{(i)}\right)_{x_{i}\neq\xi_{i}}$ are unitary operators. The unitary operators can be further fixed without loss of generality to be the identity. Explicitly, $\mathcal{S}$ being an arbitrary quantum process, the probability distribution $\mathcal{S}\left(\mathcal{M}^{(1)}_{a_{1}\mid x_{1}},\cdots,\mathcal{M}^{(N)}_{a_{N}\mid x_{N}}\right)$ can be reproduced by
$\mathcal{S}^{\prime}\left(\mathcal{M}^{(1)^{\prime}}_{a_{1}\mid x_{1}},\cdots,\mathcal{M}^{(N)^{\prime}}_{a_{N}\mid x_{N}}\right)$
$=\mathcal{S}\left(\mathcal{U}^{(1)}\circ\mathcal{M}^{(1)^{\prime}}_{a_{1}\mid x_{1}},\cdots,\mathcal{U}^{(N)}\circ\mathcal{M}^{(N)^{\prime}}_{a_{N}\mid x_{N}}\right)\,,$ (102)
where $\mathcal{M}^{(i)^{\prime}}_{a_{i}\mid x_{i}}$ are the instruments obtained from $\mathcal{M}^{(i)}_{a_{i}\mid x_{i}}$ by letting $U_{x_{i}}^{(i)}=I$, $\mathcal{U}^{(i)}$ is the controlled unitary gate $I\otimes|\xi_{i}\rangle\langle\xi_{i}|+\sum_{x_{i}\neq\xi_{i}}U_{x_{i}}^{(i)}\otimes|x_{i}\rangle\langle x_{i}|$, and the process $\mathcal{S}^{\prime}$ is the composite of the process $\mathcal{S}$ and the local operations $\left(\mathcal{U}^{(i)}\right)_{i=1}^{N}$.
The equal-rank projectors $\left(P_{a_{i},\xi_{i}}^{(i)}\right)_{a_{i}=0}^{m_{i}-1}$ can be written as the tensor product of rank-one projectors and the identity on an additional system $R^{(i)}$, i.e.
$P_{a_{i},\xi_{i}}^{(i)}=|a_{i}\rangle\langle a_{i}|\otimes I_{R^{(i)}}\,.$ (103)
Putting Eqs. (102) and (103) together, the probability distribution $\mathcal{S}\left(\mathcal{M}^{(1)}_{a_{1}\mid x_{1}},\cdots,\mathcal{M}^{(N)}_{a_{N}\mid x_{N}}\right)$ can be written as
$\mathcal{S}^{\prime\prime}\left(\mathcal{M}^{(1)*}_{a_{1}\mid x_{1}},\cdots,\mathcal{M}^{(N)*}_{a_{N}\mid x_{N}}\right)$
$=\mathcal{S}^{\prime}\left(\mathcal{M}^{(1)*}_{a_{1}\mid x_{1}}\otimes\mathcal{I}_{R^{(1)}},\cdots,\mathcal{M}^{(N)*}_{a_{N}\mid x_{N}}\otimes\mathcal{I}_{R^{(N)}}\right)\,,$ (104)
where $\mathcal{M}_{\,a_{i}|x_{i}}^{(i)*}$ are the canonical instruments, and the process $\mathcal{S}^{\prime\prime}$ is obtained by incorporating identity gates on the additional systems $R^{(i)}$ into the process $S^{\prime}$. In other words, the maximum quantum ICO value of a single-trigger correlation can be obtained by fixing the local operation of each party to be the canonical instrument $\mathcal{M}_{\,a_{i}|x_{i}}^{(i)*}$. This completes the proof.